### v2

# Large Model Based Agents: State-of-the-Art, Cooperation Paradigms, Security and Privacy, and Future Trends

Abstract
With the rapid advancement of large models (LMs), the development of general-purpose intelligent agents powered by LMs has become a reality. It is foreseeable that in the near future, LM-driven general AI agents will serve as essential tools in production tasks, capable of autonomous communication and collaboration without human intervention. This paper investigates scenarios involving the autonomous collaboration of future LM agents. We review the current state of LM agents, the key technologies enabling LM agent collaboration, and the security and privacy challenges they face during cooperative operations. To this end, we first explore the foundational principles of LM agents, including their general architecture, key components, enabling technologies, and modern applications. We then discuss practical collaboration paradigms from data, computation, and knowledge perspectives to achieve connected intelligence among LM agents. After that, we analyze the security vulnerabilities and privacy risks associated with LM agents, particularly in multi-agent settings, examining underlying mechanisms and reviewing current and potential countermeasures. Lastly, we propose future research directions for building robust and secure LM agent ecosystems.
Index Terms:
Large models, AI agents, embodied intelligence, networking, multi-agent collaboration, security, privacy.
I Introduction
I-A Background of Large Model Based Agents
In the 1950s, Alan Turing introduced the Turing Test to assess whether machines could exhibit intelligence comparable to humans. These artificial entities, commonly known as “agents”, serve as the core components of artificial intelligence (AI) systems. AI agents (also known as agentic AI [1]) are autonomous entities capable of understanding and responding to human inputs, perceiving their environment, making decisions, and taking actions across physical, virtual, or mixed-reality settings to achieve specific goals [2]. They can be software-based or physical entities, functioning independently or in collaboration with humans or other agents. Since the mid-20th century, significant progress has been made in the development of AI agents [3, 4], such as Deep Blue, AlphaGo, and AlphaZero, as shown in Fig. 1. Despite these advances, prior research primarily concentrated on refining specialized abilities such as symbolic reasoning or excelling in certain tasks such as Go or Chess, often neglecting the cultivation of general-purpose capabilities within AI models such as long-term planning, multi-task generalization, and knowledge retention.
Figure 1: Evolution history of AI agents.
With the rise of large models (LMs) [5], including large language models (LLMs) such as OpenAI GPT-4, Google PaLM 2, and Microsoft Copilot, as well as large vision models (LVMs) such as DALL-E 2, LM-based agents (also known as LM agents or agentic LMs [6]) are unlocking new possibilities [7]. LM agents significantly enhance the inherent capabilities of AI systems, providing a versatile foundation for the next-generation AI agents [8]. Serving as the “brain” of AI agents, LMs empower them with advanced capabilities in human-machine interaction (HMI), few/zero-shot planning, contextual understanding, knowledge retention, and general-purpose task solving across physical, virtual, or mixed-reality environments [9, 10, 11]. LM agents generally fall into two categories:
  • Software-form LM agents, such as AutoGPT [12] and AutoGen [13], can autonomously interpret human instructions and use various tools (e.g., search engines and external APIs) to gather information and complete intricate tasks [14]. For instance, as shown in Fig. 2(a), an LM-powered virtual assistant can generate personalized travel plans, set reminders, and manage tasks while continuously learning and adapting in dynamic environments.
  • Embodied LM agents, such as FigureAI’s Figure 02 and Tesla’s Optimus, engage directly with the physical world. These agents perceive and interact with their surroundings, allowing them to solve real-world problems [15]. For instance, as shown in Fig. 2(b), an LM-powered household robot analyzes room layouts, surface types, and obstacles, to devise customized cleaning strategies rather than merely following generic instructions.
LM agents are recognized as a significant step towards achieving artificial general intelligence (AGI) and have been widely applied across fields such as web search [16], recommendation systems [17], virtual assistants [18, 19], Metaverse gaming [20], robotics [15], autonomous vehicles [21], and automated attack penetration [22]. As reported by MarketsandMarkets [23], the worldwide market for autonomous AI and autonomous agents was valued at USD 4.8 billion in 2023 and is projected to grow at a CAGR of 43%, reaching USD 28.5 billion by 2028. LM agents have attracted global attention, and leading technology giants including Google, OpenAI, Microsoft, IBM, AWS, Oracle, NVIDIA, and Baidu are venturing into the LM agent industry.
Figure 2: Use cases of LM agents: (a) a software-form LM agent acting as the virtual assistant; (b) an embodied LM agent serving as the home cleaner. An LM agent, either in software or embodied form, generally consists of four key components: interaction, planning, action, and memory.
Figure 3: Overview of LM agent networks. Each LM agent includes two parts: (i) the brain located in the cyberspace powered by LMs such as GPT-4o, PaLM 2, and Copilot; and (ii) the physical or digital body such as autonomous vehicle, robot dog, UAV, and digital human. Within each LM agent, the cyber brain synchronizes with its physical/digital body via intra-agent communications. LM agents communicate with one another in the cloud to share task-oriented, knowledge, and inference results via inter-agent communications, establishing a network of interconnected intelligence. Each LM agent dynamically interact with other agents, virtual/real environments, and humans. The brain of each LM agent can be deployed either as a standalone entity or in a hierarchical manner across various platforms such as cloud servers, edge servers, or end devices.
I-B Motivation of LM Agent Networks
Beyond the intelligence of single LM agents, connected LM agents can form LM agent networks to cooperatively address complex tasks beyond the capacity of individual agents [24]. As depicted in Fig. 3, within LM agent networks, connected LM agents can freely share sensory data, task-oriented knowledge, and LM inference results. By leveraging collective computational power and shared expertise from specialized LM agents, it fosters cooperative decision-making and collective intelligence. For instance, in autonomous driving, connected autonomous vehicles, acting as LM agents, share real-time sensory data, coordinate movements, and negotiate passage at intersections to optimize traffic flow and enhance safety. Each LM agent consists of an LM-powered brain (hosted in the cloud, edge, or end devices) and a physical or digital body. For embodied LM agents, a tiny local LM within the physical body handles local inference, while compute-intensive tasks are offloaded to the cloud LM, enabling collaborative computation between the cloud brain and the local brain. Each LM agent can interact with other agents, virtual/real environments, and humans. Communication within LM agent networks occurs via two primary modes: (i) intra-agent communications for seamless synchronization of status, data, and knowledge between the brain and body to enable real-time functionality; and (ii) inter-agent communications for efficient information exchange and computational coordination among multiple LM agents.
LM agent networks represent a transformative leap in intelligent systems, which incorporate a range of cutting-edge technologies as its foundation. Particularly, foundation models enable cognitive capabilities through advanced reasoning and planning, knowledge-related technologies integrate external knowledge sources for context-aware actions, HMI technologies ensure seamless human-agent interactions via natural language processing (NLP) and augmented/virtual/mixed reality (AR/VR/MR), digital twin technologies synchronize physical and digital states through intra-agent communications, and multi-agent collaboration technologies foster efficient teamwork via inter-agent communications. However, their full potential faces key challenges, including but not limited:
  • Dynamic Construction of LM Agent Networks. Constructing dynamic LM agent networks requires addressing the versatility-efficiency-portability trilemma [25]. The network should support diverse tasks and applications (versatility), achieve high resource utilization with low cost (efficiency), and operate across heterogeneous platforms and environments (portability). The inherent spatiotemporal dynamics introduce additional complexity. From temporal perspectives, agents’ roles and behaviors evolve over time, necessitating adaptive reconfiguration mechanisms to ensure continuity in long-term operations. From spatial perspectives, the topologies of LM agent networks change due to mobility and environmental changes, requiring robust coordination protocols to maintain network coherence.
  • Collaborative LM Service Provisioning in Heterogeneous Networks. Once constructed, LM agent networks should enable heterogeneous agents to collaboratively provide LM services. Given the substantial computational resources required to run a complete LM, edge or terminal agents often lack the capacity to operate a full-scale LM. There exist two primary strategies: (i) LM lightweighting, e.g., knowledge distillation [26], quantization [27], pruning [28], and hardware acceleration [29], for deploying compact LM variants tailored to individual agents’ computational capacities. Collaboration is facilitated through methods such as role-playing for task-specific role assignment and distributed consensus for coordinated decision-making. (ii) LM sharding, e.g., split learning (SL) [30] and mixture-of-experts (MoE) [31], for distributed LM reconstruction, enabling agents to share workloads efficiently. The full-scale LM is segmented into smaller, manageable shards distributed across agents, allowing each edge or terminal agent to handle a portion of the model that aligns with its resource capabilities.
  • Autonomous Optimization and Secure Collaboration. Dynamic resources (e.g., data, knowledge, and computational power) allocation is vital to maximize throughput, minimize latency, and maintain seamless communication across cloud-edge-end layers. Security mechanisms in collaboration are also crucial to ensure trustworthiness (i.e., LM agents operate securely and transparently in collaborative tasks) and privacy preservation (i.e., safeguarding sensitive data during agent interactions).
I-C Motivation of Securing LM Agent Networks
Despite the promising future of connected LM agents, security and privacy concerns pose significant barriers to their widespread adoption. Throughout the life-cycle of LM agents, numerous vulnerabilities can emerge, ranging from adversarial examples [32], agent poisoning [33], LM hallucination [34], to pervasive data collection and memorization risks [35].
1) Security vulnerabilities. LM agents are susceptible to “hallucinations”, where their LMs generate outputs that are plausible but incorrect and not grounded in reality, potentially spreading misinformation in multi-agent environments [34]. Hallucinations can compromise the reliability of decision-making, cause task failures, and pose risks to both physical entities and humans. Ensuring the integrity of intermediate inference results in collaborative systems is crucial, as biased or manipulated inputs can lead to unfair or erroneous outcomes [36]. Attacks such as adversarial manipulations [32], poisoning [37], and backdoors [38] exacerbate these vulnerabilities by allowing malicious actors to manipulate inputs and deceive the models. In collaborative environments, agent poisoning behaviors [33], where malicious agents disrupt the operation of others, can undermine the collaborative systems. Additionally, integrating LM agents into cyber-physical-social systems (CPSS) expands the attack surface, providing more opportunities for malicious exploitation within interconnected systems.
2) Privacy breaches. The extensive data collection processes and data memorization capabilities of LM agents raise significant risks related to data breaches and unauthorized access. These agents often process large volumes of personal and sensitive business information in both to-customer (ToC) and to-business (ToB) applications, which raises concerns regarding data storage, processing, and sharing [39]. Furthermore, LMs can inadvertently memorize sensitive details from training data and past interactions, potentially exposing private information during subsequent interactions [35]. These privacy risks are exacerbated in multi-agent collaborations, where LM agents may unintentionally leak sensitive information about users, other agents, or their internal operations during communication and task execution.
I-D Related Surveys and Our Contributions
Recently, LM agents have garnered significant interest across academia and industry, leading to a variety of research exploring their potential from multiple perspectives. Notable survey papers in this field are as below. Andreas et al. [40] present a toy experiment for AI agent construction and case studies on modeling communicative intentions, beliefs, and desires. Wang et al. [41] identify key components of LLM-based autonomous agents (i.e., profile, memory, planning, and action) and the subjective and objective evaluation metrics. Besides, they discuss the applications of LLM agents in engineering, natural science, and social science. Xi et al. [10] present a general framework for LLM agents consisting of brain, action, and perception. Besides, they explore applications in single-agent, multi-agent, and human-agent collaborations, as well as agent societies. Zhao et al. [42] offer a systematic review of LLMs in terms of pre-training, adaptation tuning, utilization, and capacity assessment. Besides, background information, mainstream technologies, and critical applications of LLMs are introduced. Xu et al. [43] provide a tutorial on key concepts, architecture, and metrics of edge-cloud AI-generated content (AIGC) services in mobile networks, and identify several use cases and implementation challenges. Huang et al. [2] offer a taxonomy of AI agents in virtual/physical environments, discuss cognitive aspects of AI Agents, and survey the applications of AI agents in robotics, healthcare, and gaming. Cheng et al. [11] review key components of LLM agents (including planning, memory, action, environment, and rethinking) and their potential applications. Planning types, multi-role relationships, and communication methods in multi-agent systems are also reviewed. Masterman et al. [9] provide an overview of single-agent and multi-agent architectures in industrial projects and present the insights and limitations of existing research. Guo et al. [44] discuss the four components (i.e., interface, profiling, communication, and capabilities acquisition) of LLM-based multi-agent systems and present two lines of applications in terms of problem solving and world simulation. Durante et al. [45] introduce multimodal LM agents and a training framework including learning, action, congnition, memory, action, and perception. They also discuss the different roles of agents (e.g., embodied, simulation, and knowledge inference), as well as the potentials and experimental results in different applications including gaming, robotics, healthcare, multimodal tasks, and NLP. Hu et al. [20] outline six key components (i.e., perception, thinking, memory, learning, action, and role-playing) of LLM-based game agents and review existing LLM-based game agents in six types of games. Xu et al. [5] provide a comprehensive survey of enabling architectures and challenges for LM agents in gaming. Qu et al. [46] provide a comprehensive survey on integrating mobile edge intelligence (MEI) with LLMs, emphasizing key applications of deploying LLMs at the network edge along with state-of-the-art techniques in edge LLM caching, delivery, training, and inference.
Existing surveys on LM agents primarily focus on the general framework design of LLM agents and multi-agent systems, as well as their potential in specific applications. Distinguished from the above-mentioned existing surveys, this survey focuses on the networking aspect of LM agents, including the general architecture, enabling technologies, key characteristics, and collaboration paradigms to construct networked systems of LM agents within physical, virtual, or mixed-reality environments. Additionally, with the advances of LM agents, it is urgent to examine their security and privacy in future AI agent systems. This work comprehensively reviews the security and privacy issues of LM agents and discusses both existing and potential defense mechanisms, which are overlooked in previous surveys. Table I compares the contributions of our survey with previous related surveys in the field of LM agents.
TABLE I: A Comparison of Our Survey with Relevant Surveys
Year. Refs. Contribution
2022 [40]
A toy experiment for AI agent construction and case studies
in modeling communicative intentions, beliefs, and desires.
2023 [41]
Survey on key compoments and evaluation policies of LLM
agents, and applications in engineering, natural science, and
social science.
2023 [10]
Discussions on general framework for LLM agents and
applications of single-agent, multi-agent, and human-agent
collaborations, as well as agent societies.
2023 [42]
Review on background, pre-training, adaptation tuning,
evaluation & utilization, capacity assessment, and critical
applications of LLMs.
2024 [43]
Tutorial on key concepts, architecture, and metrics of
edge-cloud AIGC services in mobile networks, and
identify use cases and key implementation challenges.
2024 [2]
Discussions on taxonomy of AI agents in virtual/physical
environments, cognitive aspects, and applications in robots,
healthcare, and gaming.
2024 [11]
Discuss key components and applications of LLM agents,
and review planning types, multi-role relationships, and
communication modes in multi-agent systems.
2024 [9]
Overview of single-agent and multi-agent architectures in
industrial projects and insights and limitations of research.
2024 [44]
Discuss key components of LLM-based multi-agent systems
and applications in problem solving and world simulation.
2024 [45]
Discuss key concepts and the framework of multimodal LM
agents, the different roles of agents, and the potentials and
experimental results in gaming, robotics, healthcare, NLP,
and multimodality.
2024 [20]
Discuss key components of LLM-based game agents and
review existing approaches in six types of games.
2024 [5]
Survey the enabling architectures and key challenges of LM
agents for games.
2024 [46]
Survey on key applications of deploying LLMs at network
edges and state-of-the-art techniques in edge LLM caching,
delivery, training, and inference.
Now Ours
Comprehensive survey of the fundamentals, security, and
privacy of connected LM agents, discussions on the general
architecture, enabling technologies, key characteristics, and
cooperation paradigms of connected LM agents, discussions
on security/privacy threats, state-of-the-art countermeasures,
and open research issues in future LM agent systems.
Figure 4: Organization structure of this paper.
In this paper, we present a systematic review of the state-of-the-arts in both single and connected LM agents, focusing on cooperation paradigms, security and privacy threats, existing and potential countermeasures, and future trends. Our survey aims to 1) provide a broader understanding of how LM agents work and how they cooperate in multi-agent scenarios, 2) examine the scope and impact of security and privacy challenges associated with LM agents and their interactions, and 3) highlight effective strategies and solutions for defending against these threats to safeguard LM agents in various intelligent applications. The main contributions of this work are four-fold.
  • We comprehensively review recent advances in LM agent construction across academia and industry. We investigate the fundamentals of LM agents, including the general architecture, and key components (i.e., planning, memory, action, interaction, and security modules). Industrial prototypes and potential applications of LM agents are also discussed.
  • We discuss key characteristics and enabling technologies of LM agent networks and explore practical cloud-edge-end collaboration paradigms from the aspects of data cooperation, computation cooperation, and knowledge cooperation. We systematically discuss interaction strategies of LM agents (i.e., cooperation, partial cooperation, and competition).
  • We comprehensively analyze existing and potential security and privacy threats, their underlying mechanisms, categorization, and challenges for both single and connected LM agents. We also review state-of-the-art countermeasures and examine their feasibility in securing LM agent networks.
  • Lastly, we discuss open research opportunities and point out future research directions of LM agent networks, aiming to inspire more ongoing research and innovations in this field.
I-E Paper Organization
The remainder of this paper is organized as below. Section II gives an overview of single LM agents, while Section III presents the fundamentals for networking LM agents. Section IV and Section V introduce the taxonomy of security and privacy threats to LM agents, respectively, along with state-of-the-art countermeasures. Section VI outlines open research issues and future directions in the field of LM agents. Finally, conclusions are drawn in Section VII. Fig. 4 depicts the organization structure of this survey.
TABLE II: Progress of Standards for LM Agents
Standard Publication Date Main Content
IEEE SA - P3394 2023-09-21
The natural language interface is defined, including various protocols and guidelines that enable applications
and agents to effectively communicate with LLM agents, such as API syntax and semantics.
IEEE SA - P3428 2023-12-06
The integration of LLMs with existing education systems to ensure that LLMs can seamlessly interact with
AIS while addressing issues of bias, transparency, and accountability in educational environments.
II An Overview of Large Model Agents
In this section, we first introduce existing standards of LM agents. Then, we discuss the operating system (OS) and constructing modules of LM agents. Next, we introduce typical prototypes and discuss modern applications of LM agents.
Figure 5: Illustration of the OS architecture of LM agents [47, 48].
II-A Standards of LM Agents
We briefly introduce two existing standards on LM agents: IEEE SA-P3394 and IEEE SA-P3428.
1) The IEEE SA - P3394 standard [^1] https://standards.ieee.org/ieee/3394/11377/, launched in 2023, defines natural language interfaces to facilitate communication between LLM applications, agents, and human users. This standard establishes a set of protocols and guidelines that enable applications and agents to effectively communicate with LLM agents. These protocols and guidelines include, but are not limited to, API syntax and semantics, voice and text formats, conversation flow, prompt engineering integration, LLM thought chain integration, as well as API endpoint configuration, authentication, and authorization for LLM plugins. The standard is expected to advance technological interoperability, promote AI industry development, enhance the practicality and efficiency of LMs, and improve AI agents’ functionality and user experience.
2) The IEEE SA - P3428 standard [^2] https://standards.ieee.org/ieee/3428/11489/, launched in 2023, aims to develop standards for LLM agents in educational applications. The primary goal is to ensure the interoperability of LLM agents across both open-source and proprietary systems. Key areas of focus include the integration of LLMs with existing educational systems and addressing technical and ethical challenges. This includes ensuring that LLMs can seamlessly interact with other AI components, such as adaptive instructional systems (AIS), while addressing issues of bias, transparency, and accountability within educational contexts. The standard is intended to support the widespread and effective application of LLMs in the field of education, thereby enabling more personalized, efficient, and ethically sound AI-driven educational experiences.
II-B Operating System (OS) of LM Agents
According to [47, 48], the OS architecture of LM agents consists of three layers: application, kernel, and hardware, as shown in Fig. 5.
  • The application layer hosts agent applications (e.g., travel, coding, and robot agents) and offers an SDK that abstracts system calls, simplifying agent development. Typical agent developing frameworks include LangChain, AutoGen, Dify, and LlamaIndex.
  • The kernel layer includes the ordinary OS kernel (to process LM-related queries such as file operation and network request) and an additional LM agent kernel (to process LM-related queries such as reasoning, planning, and tool parsing), with a focus on without altering the original OS structure. Key modules in the LM agent kernel [47, 48] include the agent scheduler for task planning and prioritization, context manager for LM status management, memory manager for short-term data, storage manager for long-term data retention, tool manager for external API interactions, and access manager for privacy controls.
  • The hardware layer comprises physical resources (CPU, GPU, memory, etc.), which are managed indirectly through OS system calls, as LM kernels do not interact directly with the hardware.
Figure 6: Workflow of the five constructing modules of LM agents in interacting with the real/virtual environment.
II-C Constructing Modules of LM Agents
According to [9, 2, 10, 11], there are generally five constructing modules of LM agents: planning, action, memory, interaction, and security. As depicted in Fig. 6, these modules together enable LM agents to perceive, plan, act, learn, and interact efficiently and securely in complex and dynamic environments. Particularly, empowered by LMs, the planning module produces strategies and action plans with the help of the memory module, enabling informed decision-making. The action module executes these embodied actions, adapting actions based on real-time environmental feedback to ensure contextually appropriate responses. The memory module serves as a repository of accumulated knowledge (e.g., past experiences and external knowledge), facilitating continuous learning and improvement. The interaction module enables effective communication and collaboration with humans, other agents, and environment. The security module is integrated throughout LM agents’ operations, ensuring active protection against threats and maintaining integrity and confidentiality of data and processes.
Figure 7: Illustration of five constructing modules (i.e., planning, action, memory, interaction, and security) of connected LM agents including their key components.
Figure 8: Illustration of typical feedback-free planning methods of LMs: (a) basic I/O; (b) chain-of-thought (CoT) [49]; (c) tree-of-thought (ToT) [50]; (d) graph-of-thought (GoT) [51].
Figure 9: Illustration of ReAct [52]: a typical feedback-enhanced planning methods of LMs.
Figure 10: Illustration of typical feedback-enhanced planning methods of LMs: (a) Reflexion [53]; (b) ground planning [54].
II-C1 Planning Module
The planning module serves as the core of an LM agent [8, 11]. It utilizes advanced reasoning techniques to devise effective solutions fo agents to complex problems. It includes the following working modes.
  • Feedback-free planning: The planning module enables LM agents to understand the complex problems and find reliable solutions by breaking them down into necessary steps or manageable sub-tasks [8, 12], as shown in Fig. 8. For instance, chain-of-thought (CoT) [49] is a popular sequential reasoning approach where each thought builds directly on the previous one. It represents the step-by-step logical thinking and can enhance the generation of coherent and contextually relevant responses. Tree-of-thought (ToT) [50] organizes reasoning as a tree-like structure, exploring multiple paths simultaneously. In ToT, each node represents a partial solution, allowing the model to branch and backtrack to find the optimal answer. Graph-of-thought (GoT) [51] models reasoning using an arbitrary graph structure, allowing more flexible information flow. GoT captures complex relationships between thoughts, enhancing the model’s problem-solving capabilities. AVIS [55] further refines the tree search process for visual QA tasks using a human-defined transition graph and enhances decision-making through a dynamic prompt manager.
  • Feedback-enhanced planning: To make effective long-term planning in complex tasks, it is necessary to iteratively reflect on and refine execution plans based on past actions and observations [41]. The goal is to correct past errors and improve final outcomes. For instance, as shown in Fig. 9, ReAct [52] combines reasoning and acting by prompting LLMs to generate reasoning traces (i.e., thoughts) and actions simultaneously. This dual approach allows the LLM to create, monitor, and adjust action plans, while task-specific actions enhance interaction with external sources, improving response accuracy and reliability. As shown in Fig. 10(a), Reflexion [53] converts environmental feedback into self-reflection and enhances ReAct by enabling LLM agents to learn from past errors and iteratively optimize behaviors. Reflexion features an actor that produces actions and text via models (e.g., CoT and ReAct) enhanced by memory, an evaluator that scores outputs using task-specific reward functions, and self-reflection that generates verbal feedback to improve the actor.
  • Multi-persona self-planning: Inspired by pretend play, Wang et al. [56] develop a cognitive synergist that enables a single LLM to split into multiple personas, facilitating self-collaboration for solving complex tasks. They propose solo performance prompting (SPP), where LLM identifies, simulates, and collaborates with diverse personas, such as domain experts or target audiences, without external retrieval systems. SPP enhances problem-solving by allowing LLM to perform multi-turn self-revision and feedback from various perspectives.
  • Grounded planning: Executing plans in real or simulated world environments (e.g., Minecraft) requires precise, multi-step reasoning. As shown in Fig. 10(b), VOYAGER [54], the first LLM-powered agent in Minecraft, utilizes in-context lifelong learning to adapt and generalize skills to new tasks and worlds. VOYAGER include an automatic curriculum for exploration, a skill library of executable code for complex behaviors, and an iterative prompting mechanism that refines programs based on feedback. Wang et al. [57] further propose an interactive describe-explain-plan-select (DEPS) planning approach that improves LLM-generated plans by integrating execution descriptions, self-explanations, and a goal selector that ranks sub-goals to refine planning.
II-C2 Memory Module
The memory module is integral to LM agent’s ability to learn and adapt over time [41]. It maintains an internal memory that accumulates knowledge from past interactions, thoughts, actions, observations, and experiences with users, other agents, and the environments. The stored information guides future decisions and actions, allowing the agent to continuously refine its knowledge and skills. This module ensures that the agent can remember and apply past lessons to new situations, thereby improving its long-term performance and adaptability [11]. There are various memory formats such as embedded vectors, databases, and structured lists. Additionally, retrieval-augmented generation (RAG) technologies [58] are employed to access external knowledge sources, further enhancing the accuracy and relevance of LM agent’s planning capabilities. In the literature [41, 11], memory can be divided into the following three types.
  • Short-term memory focuses on the contextual information of the current situation. It is temporary and limited, typically managed through a context window that restricts the amount of information the LM agent can learn at a time.
  • Long-term memory stores LM agent’s historical behaviors and thoughts. This is achieved through external vector storage, which allows for quick retrieval of important information, ensuring that the agent can access relevant past experiences to inform current decisions [59].
  • Hybrid memory combines short-term and long-term memory to enhance an agent’s understanding of the current context and leverage past experiences for better long-term reasoning. Liu et al. [60] propose the RAISE architecture to enhance ReAct for conversational AI agents by integrating a dual-component memory system, where Scratchpad captures recent interactions as short-term memory; while the retrieval module acts as long-term memory to access relevant examples.
II-C3 Action Module
The action module equips the LM agent with the ability to execute and adapt actions in various environments [10, 45]. This module is designed to handle embodied actions and tool-use capabilities, enabling the agent to interact with its physical surroundings adaptively and effectively. Besides, tools significantly broaden the action space of the agent.
  • Embodied actions. The action module empowers LM agents to perform contextually appropriate embodied actions and adapt to environmental changes, facilitating interaction with and adjustment to physical surroundings [61, 15]. As LLM-generated action plans are often not directly executable in interactive environments, Huang et al. [61] propose refining LLM-generated plans for embodied agents by conditioning on demonstrations and semantically translating them into admissible actions. Evaluations in the VirtualHome environment show significant improvements in executability, ranging from 18% to 79% over the baseline LLM. Besides, SayCan [15] enables embodied agents such as robots to follow high-level instructions by leveraging LLM knowledge in physically grounded tasks, where LLM (i.e., Say) suggests useful actions; while learned affordance functions (i.e., Can) assess feasibility. SayCan’s effectiveness is demonstrated through 101 zero-shot real-world robotic tasks in a kitchen setting. PaLM-E [62] is a versatile multimodal language model for embodied reasoning, visual-language, and language tasks. It integrates continuous sensor inputs, e.g., images and state estimates, into the same embedding space as language tokens, allowing for grounded inferences in real-world sequential decision-making.
  • Learning to use & make tools. By leveraging various tools (e.g., search engines and external APIs) [63], LM agents can gather valuable information to handle assigned complex tasks. For instance, AutoGPT integrates LLMs with predetermined tools such as web and file browsing. InteRecAgent [17] integrates LLMs as the brain and recommender models as tools, using querying, retrieval, and ranking tools to handle complex user inquiries. Beyond using existing tools, LM agents can also develop new tools to enhance task efficiency [10]. To optimize tool selection with a large toolset, ReInvoke [14] introduces an unsupervised tool retrieval method featuring a query generator to enrich tool documents in offline indexing and an intent extractor to identify tool-related intents from user queries in online inference, followed by a multi-view similarity ranking strategy to identify the most relevant tools.
II-C4 Interaction Module
The interaction module enables the LM agent to interact with humans, other agents, and the environment [44]. Through these varied interactions, LM agents can gather diverse experiences and knowledge, which are essential for comprehensive understanding and adaptation.
  • Agent-Agent Interactions. LM agents can coordinate efforts on shared tasks, exchange knowledge to solve problems, and negotiate roles to collaborate with other agents, fostering a cooperative network [63]. Theory of mind [64] can enhance the ability to understand other agents’ hidden mental states. Li et al. [64] propose a prompt-engineering method to incorporate explicit belief state representations. They also introduce a novel evaluation of LLMs’ high-order theory of mind in teamwork scenarios, emphasizing dynamic belief state evolution and intent communication among agents.
  • Agent-Human Interactions. LM agents can interact with humans by understanding natural languages, recognizing human emotions and expressions, and providing assistance in various tasks [20]. As found by work [65], LLMs such as GPT-4 tend to forget character settings in multi-turn dialogues and struggle with detailed role assignments due to context window limits. To address this, a tree-structured persona model is devised in [65] for character assignment, detection, and maintenance, enhancing agent interactions.
  • Agent-Environment Interactions. LM agents can engage directly with the physical or virtual environments. By facilitating engagement in physical, virtual, or mixed-reality environments [15, 2], the interaction module ensures that LM agents can operate effectively across different contexts. Lai et al. develop the AutoWebGLM agent [66], which excels in web browsing tasks through curriculum learning, self-sampling reinforcement learning, and rejection sampling fine-tuning. A Chrome extension based on AutoWebGLM validates its effective reasoning and operation capability across various websites in real-world services.
II-C5 Security Module
The security module is crucial to ensure the secure, safe, ethical, and privacy-preserving operations of LM agents [45]. It is designed to monitor and regulate the LM agents’ actions, interactions, and decisions to prevent harm and ensure compliance with legal and ethical standards. LMs can be utilized to design advanced and cost-effective defense mechanisms, such as PentestGPT [22] for automated network vulnerability detection. LMs also introduce new threats to agents such as magnified hallucination and memorization risks. The security module should dynamically adapt to emerging threats by learning from new security/privacy incidents and integrating updates from security/privacy databases and policies. It also incorporates ethical guidelines and bias mitigation techniques to ensure fair and responsible behaviors.
Connections Between Modules: The above five key components of an LM agent are interconnected to create a cohesive and intelligent system. Particularly, the planning module relies on the memory module to access past experiences and external knowledge, ensuring informed decision-making. The action module executes plans generated by the planning module, adapting actions based on real-time feedback and memory. The interaction module enhances these processes by facilitating communication and collaboration, which provides additional data and context for the planning and memory modules. Besides, security considerations are seamlessly integrated into every aspect of the LM agent’s operations to ensure robust and trusted executions.
TABLE III: Comparison of Existing Typical LM Agent Prototypes
Prototype Application
Enhancing
Intelligence
Improving
Safety
Optimizing
Experience
Other Key Features
AutoGPT General scenarios Yes N/A Yes Task decomposition and execution
AutoGen General scenarios Yes N/A Yes Multi-agent conversation framework
BabyAGI General scenarios Yes N/A Yes Autonomous task management simulation
ChatDev General scenarios Yes N/A Yes Conversational AI for software development
MetaGPT General scenarios Yes N/A Yes Meta-learning for task adaptation
MobileAgent v2 General scenarios Yes N/A Yes Multimodal input support
AppAgent General scenarios Yes N/A Yes Performs tasks on mobile devices
NetLLM Mobile communications Yes N/A Yes Fine-tunes LLM for networking
NetGPT Mobile communications Yes N/A Yes Cloud-edge cooperative LM service
Figure 02 by FigureAI Intelligent robots Yes Yes N/A Performs dangerous jobs
Optimus by Tesla Intelligent robots Yes Yes N/A Second-generation humanoid robot
Baidu Apollo ADFM Autonomous driving Yes Yes Yes Supports L4 autonomous driving
PentestGPT Attack-defense confrontation Yes Yes N/A 87% success in vulnerability exploitation
AutoAttacker Attack-defense confrontation Yes No N/A Automatically execute network attacks
II-D Modern Prototypes & Applications of LM Agents
II-D1 General Scenarios
Recently, various industrial projects of LM agents, such as AutoGPT, AutoGen, BabyAGI, ChatDev, and MetaGPT, demonstrate their diverse potential in assisting web, life, and business scenarios, such as planning personalized travels, automating creative content generation, and enhancing software development workflows. For instance, AutoGPT [^3] https://autogpt.net/ is an open-source autonomous agent utilizing GPT-3.5 or GPT-4 APIs to independently execute complex tasks by breaking down them into several sub-tasks and chaining LLM outputs, showcasing advanced reasoning capabilities [12]. AutoGen [^4] https://microsoft.github.io/autogen/, developed by Microsoft, offers an open-source multi-agent conversation framework, supports APIs as tools for improved LLM inference, and emphasizes the automatic generation and fine-tuning of AI models [13]. BabyAGI [^5] https://github.com/yoheinakajima/babyagi integrates task management via OpenAI platforms and vector databases, simulating a simplified AGI by autonomously creating and executing tasks based on high-level objectives. ChatDev [^6] https://github.com/OpenBMB/ChatDev focuses on enhancing conversational AI, providing sophisticated dialogue management, coding, debugging, and project management to streamline software development processes [67]. MetaGPT [^7] https://www.deepwisdom.ai/ explores the meta-learning paradigm, where the model is trained to rapidly adapt to new tasks by leveraging knowledge from related tasks, thus improving performance across diverse applications [68].
LM agents can also enhance the QoE of end users. For instance, MobileAgent v2 [18], launched by Alibaba, is a mobile device operation assistant that achieves effective navigation through multi-agent collaboration, automatically performing tasks such as map navigation, and supports multimodal input including visual perception, enhancing operational efficiency on mobile devices. AppAgent [19], developed by Tencent, performs various tasks on mobile phones through autonomous learning and imitating human click and swipe gestures, including posting on social media, helping users write and send emails, using maps, online shopping, and even complex image editing.
II-D2 Mobile Communications
LM agents offer significant advantages for mobile communications by enabling low-cost and context-aware decision-making [69], personalized user experiences [70], and automatic optimization problem formulation for wireless resource allocation [71]. For instance, NetLLM [69] fine-tunes the LLM to acquire domain knowledge from multimodal data in networking scenarios (e.g., adaptive bitrate streaming, viewport prediction, and cluster job scheduling) with reduced handcraft costs. Meanwhile, NetGPT [70] uses a cloud-edge cooperative LM framework for personalized outputs and enhanced prompt responses in mobile communications via de-duplication and prompt enhancement technologies. ChatNet [72] uses four GPT-4 models to serve as analyzer (to plan network capacity and designate tools), planner (to decouple network tasks), calculator (to compute and optimize the cost), and executor (to produce customized network capacity solutions) via prompt engineering.
II-D3 Intelligent Robots
LM agents play a crucial role in advancing intelligent industrial and service robots [15]. These robots can perform complex tasks such as product assembly, environmental cleaning, and customer service, by perceiving surroundings and learning necessary skills through LMs. In August 2024, FigureAI released Figure 02888https://spectrum.ieee.org/figure-new-humanoid-robot, a human-like robot powered by OpenAI LM, capable of fast common-sense visual reasoning and speech-to-speech conversation with humans to handle dangerous jobs in various environments. Besides, in July 2024, Tesla unveiled its second-generation humanoid robot named Optimus [^9] https://viso.ai/edge-ai/tesla-bot-optimus/, which demonstrates the enhanced capabilities brought by LM agents.
II-D4 Autonomous Driving
LM agents are transforming autonomous driving by enhancing vehicle intelligence, improving safety, and optimizing driving experience (e.g., offering personalized in-car experience) [21]. In May 2024, Baidu’s Apollo Day saw the launch of Carrot Run’s sixth-generation unmanned vehicle, which is built upon the Apollo autonomous driving foundation model (ADFM) [^10] https://autonews.gasgoo.com/icv/70033042.html, the first LM agent supporting L4 autonomous driving. Companies such as Tesla, Waymo, and Cruise are also leveraging LM agents to refine their autonomous driving systems, aiming for safer and more efficient transportation systems.
Figure 11: General architecture of LM agent networks in bridging the human, physical, and cyber worlds. The LM agent has five constructing modules: planning, action, memory, interaction, and security modules. The engine of connected LM agents is empowered by a combination of five cutting-edge technologies: foundation models, knowledge-related technologies, interaction, digital twin, and multi-agent collaboration. LM agents interact with humans through HMI technologies such as NLP, with the assistance of handheld devices and wearables, to understand human intentions, desires, and beliefs. LM agents can synchronize data and statuses between the physical body and the digital brain through digital twin technologies, as well as perceive and act upon the surrounding virtual/real environment. LM agents can be interconnected in the cyber space through efficient cloud-edge networking for efficient sharing of data, knowledge, and computation results to promote multi-agent collaboration.
II-D5 Autonomous attack-defense confrontation
[…]
II-E Summary and Lessons Learned
This section provides a comprehensive review of the design, functionality, and applications of LM agents, focusing on their architectural standards, operational frameworks, and potential use cases. We identify five core modules, i.e., planning, action, memory, interaction, and security, as essential components for enabling LM agents to operate effectively in dynamic and complex environments. LM agents are already making a transformative impact across industries such as education, mobile communications, robotics, autonomous driving, and cybersecurity. To summarize, the key lessons learned include:
  • Beyond general-purpose systems like Windows and Linux, OS for embodied LM agents should cater to robotic environments such as UAVs, autonomous vehicles, and robots. Integrating systems such as ROS (Robot Operating System) for machine-centric applications will be key to ensure seamless functionality and promote embodied intelligence.
  • LMs significantly enhance LM agents’ abilities across core modules including planning, action, memory, interaction, and security, facilitating more intelligent, adaptive decision-making and improving overall performance.
  • As key enablers of next-generation AI systems, LM agents will drive the future of intelligent systems. In everyday life (e.g., home automation, transportation, personal assistance), they will empower smarter, more efficient living. In the future network landscape, LM agents will be central to evolving network dynamics, with cybersecurity shifting to agent-versus-agent defense and attack strategies. As the popularity of LM agents, efficiently constructing LM agent networks to foster collaboration and achieve synergistic effects becomes increasingly crucial.
TABLE IV: A Summary of LLM Stages, Utilized Technologies, and Representives
Stage of LLM Description Utilized Technology Representive
Required
Dataset Scale
Required
Computation Power
Pre-training
Pre-train LLM in a self-supervised
manner on a large corpus
Transformer
GPT-3, PaLM-2,
LLaMA-2
Internet-scale (e.g.,
2T tokens in LLaMA-2)
Very Large (e.g., Thousands of
GPUs for months in LLaMA-2)
Fine-tuning
Fine-tuning pretrained
LLM for downstream tasks
Instruction-tuning,
Alignment-tuning,
Transfer learning
WebGPT[16], T0,
LLaMA-2-Chat
Domain-scale (e.g.,
[…]
in LLaMA-2)
Medium (e.g., Dozens of
GPUs for days in LLaMA-2)
Prompting
Setup prompts and query trained
LLMs for generating responses
Prompt engineering,
Zero-shot prompting,
In-context learning
RALM[75],
ToolkenGPT[76]
Small-scale
Low
III Large Model Agent Networks: Working Principles
In this section, we first explore the general architecture of LM agent networks, covering their definition and underlying engine. We then discuss the key characteristics and communication modes of LM agent networks. Next, we introduce the cloud-edge-end cooperative architecture for networking LM agents, followed by an in-depth analysis of interaction strategies, including cooperation, partial cooperation, and competition. Finally, we examine cooperation paradigms in LM agent networks, focusing on data, computation, and knowledge sharing.
III-A General Architecture of LM Agent Networks
III-A1 LM Agent Networks
LM agent networks are decentralized networks of connected LM agents (either in embodied or software forms) that collaborate to accomplish complex tasks by seamlessly sharing data, knowledge, and resources. Each agent, equipped with specialized capabilities, is assigned specific sub-tasks, and collectively, these agents combine their outputs to provide comprehensive solutions. In the context of cloud-edge computing, LM agent networks optimize resource utilization and task distribution by deploying smaller, specialized models on edge devices. This ensures low-latency and scalable inference services across a range of applications, while leveraging the cloud for more resource-intensive operations. These agents can implement various collaboration strategies, such as horizontal, vertical, or hybrid models (as detailed in Sect. III-F), to enhance task execution efficiency and system performance.
III-A2 Engine of LM Agent Networks
The LM agent networks are powered by a combination of cutting-edge technologies including foundation models, knowledge-related technologies, interaction, digital twin, and multi-agent collaboration.
  • Foundation models, such as LLMs, LVMs, and large multimodal models (LMMs), serve as the brains of LM agents that comprehensively power them in planning, action, memory, interaction, and security capacities, as detailed in Sect. II-C. Table IV summarizes the basic stages of LLM services. (i) Advanced reasoning: LM technologies empower AI agents with advanced reasoning abilities such as multi-persona self-planning and grounded planning. For instance, CoT [49] and ToT [50] reasoning allow LM agents to break down complex tasks into manageable sub-tasks. (ii) Few/zero-shot generalization: Pretrained on extensive corpora, LMs demonstrate few-shot and zero-shot generalization capabilities [61], allowing LM agents to transfer knowledge seamlessly between tasks. (iii) Tool usage capability: Through the use of various tools, LM agents can gather and integrate valuable information from diverse sources to execute and adapt actions under complex scenarios. (iv) Adaptability: Through continuous learning and adaptation, LM agents can accumulate knowledge over time by learning from new data and experiences [77].
  • Knowledge-related technologies enhance LM agents by incorporating both internal knowledge (arisen from agent’s interactions with humans, environment, and other agents) and external knowledge sources (e.g., knowledge graph (KG) and vector database [78]) to produce up-to-date and contextually relevant outputs. (i) Knowledge sharing: To accomplish a common task, the locally ongoing updated private knowledge or experience of LM agents should be synchronized across all collaborators, enabling them to incrementally incorporate new information and adapt to evolving environments. (ii) Knowledge fusion: It enables LM agents to build comprehensive knowledge bases by integrating knowledge from diverse sources [79], such as structured databases, vector databases, and KGs. (iii) Knowledge retrieval: RAG combines retrieval mechanisms with generative models, enabling LM agents to dynamically fetch relevant external information from vast knowledge sources [80], which ensures the outputs of LM agents are reliable and up-to-date.
  • Interaction technologies enhance LM agents’ ability to engage naturally, immersively, and contextually with users [45]. (i) HMI: HMI technologies [15] allow LM agents to understand complex instructions, recognize speech, and interpret emotions, facilitating intuitive interactions. Besides, LM-empowered multimodal interfaces enable responses to various inputs (e.g., text, speech, and gestures), making interactions flexible and user-friendly. (ii) 3D digital humans: 3D digital humans provide realistic interfaces for empathetic and personable interactions in applications such as customer support and healthcare [43]. (iii) AR/VR/MR: AR, VR, and MR technologies create immersive and interactive environments, allowing users to engage with digital and physical elements seamlessly.
  • Digital twin technologies allow efficient and seamless synchronization of data/statuses between the physical body and the digital brain of an LM agent [81]. (i) Virtual-physical synchronization: LM agents can seamlessly synchronize attributes, behaviors, states, and other data between their bodies and brains via intra-agent bidirectional communications. The virtual representations of LM agents’ bodies can be continuously updated with real-time data inputs [81]. (ii) Virtual-physical feedback: This continuous feedback loop enhances LM agent’s contextual awareness, allowing for immediate adjustments to changing conditions. (iii) Predictive analytics: Digital twins facilitate predictive analytics and simulation, allowing LM agents to anticipate future states and optimize actions accordingly [21].
  • Multi-agent collaboration technologies enable coordinated efforts of multiple LM agents to achieve common goals and tackle complex tasks [44]. Typical technologies include multi-agent reinforcement learning (MARL) [82], cooperative game [83], mean-field game, Nash bargaining, and swarm intelligence algorithms. Specifically, collaborative problem-solving techniques, such as multi-agent planning [84] and distributed reasoning [85], enable LM agents to jointly analyze complex issues, devise solutions, and coordinate actions. Besides, efficient synchronization of knowledge allows LM agents to build upon each other’s experiences, accelerating learning in accomplishing the long-term collective task[77].
TABLE V: A Summary of Intra-Agent and Inter-Agent Communications for Connected LM Agents
Intra-agent Comm. Inter-agent Comm.
Involved Entity
[…]
Among planning, action, memory,
interaction, and security modules
[…]
Connection Type Within a single LM agent Among multiple LM agents
Support Two-way Communication
Support Multimodal Interaction
Support Semantic-aware Communication
Typical Communication Environment Wired/Wireless Wired/Wireless
III-B Key Characteristics of LM Agent Networks
LM agent networks exhibit the following distinct features that enable scalable, intelligent, and adaptive multi-agent collaboration across diverse applications.
1) High Heterogeneity. LM agent networks generally manage a vast and diverse array of nodes to support a wide range of tasks and services. Such heterogeneity also entails inherent interoperability challenges. The heterogeneity spans:
  • Capability variance: LM agents exhibit significant diversity in computational power, memory, and GPU capacities, requiring the network to effectively accommodate agents with varying resource constraints.
  • Service diversity: The LM agent network should adapt to diverse service scenarios with varying numbers of participants to ensure scalability.
  • Communication heterogeneity: LM agents generally employ diverse communication protocols and interfaces, such as cellular, satellite, and WiFi.
2) Spatiotemporal Dynamics. LM agent networks evolve dynamically in both temporal and spatial dimensions [86]. These dynamic properties require the system’s resilience and flexibility in response to real-time changes.
  • Temporal dynamics: Agents’ decision-making and interactions follow a temporal sequence, adapting to the progression of tasks or changes in the operational environment.
  • Spatial dynamics: The network topology of agents evolves due to mobility or environmental factors.
3) Semantic-aware Communication. Unlike traditional data-centric communication paradigms, LM agents leverage semantic-aware communication, prioritizing the exchange of meaningful or contextually relevant information for reduced communication overheads. Semantic-aware intra-agent or inter-agent communications facilitate a richer contextual understanding of the environment with improved accuracy of predictions and actions.
  • Intra-agent Semantic-aware Communication. Intra-agent communication focuses on synchronizing the brain and body of LM agents. These communications is semantic-aware, meaning only contextually relevant data such as changes in the environment are transmitted to minimize bandwidth usage. For instance, an LM-powered household robot sends its brain in the cloud only the changes in the environment and its position, rather than transmitting the entire environmental data repeatedly.
  • Inter-agent Semantic-aware Communication. Inter-agent communication between LM agents is inherently semantic-aware, with agents exchanging information based on relevance to their tasks or objectives. For instance, LM agents communicate via natural language for effective task coordination, negotiation, role allocation, and sequencing during task execution, and reach consensus on the final outcome.
4) Information-centric Routing. LM agent networks prioritize quickly retrieving relevant information over specific data sources. Unlike IP-based, host-oriented networks, information-centric routing focuses on “what” data is needed, enabling flexible and scalable communication through paradigms such as named data networking (NDN) and publish/subscribe (pub/sub).
  • In NDN, agents request data by content name via interest messages, retrieving it from the nearest cache to reduce latency and network traffic, with in-network caching further enhancing efficiency by storing frequently requested data locally.
  • The pub/sub model routes requests through distributed hash tables (DHTs), allowing publishers to publish content to multiple topics while subscribers continuously receive updates from their subscribed topics. For instance, MetaGPT [68] incorporates pub/sub mechanisms in multi-agent communication, enabling LLM agents to exchange messages seamlessly via a shared pool, where they publish outputs and access others’ content transparently.
5) Hierarchically Distributed Collaboration. LM agent networks employ a hierarchical yet collaborative decision-making framework across cloud, edge, and end layers. This distributed framework enables scalable and effective decision-making, particularly in resource-constrained or latency-sensitive scenarios.
  • Intra-layer collaboration: Agents within each layer (cloud, edge, or end) autonomously optimize local decisions and resource allocations.
  • Cross-layer coordination: The network seamlessly integrates cloud-scale processing, edge-level real-time responsiveness, and end-level contextual awareness to support distributed collaboration across different layers.
III-C Communication Mode of LM Agents
Every LM agent consists of two parts: (i) the LM-empowered cyber brain located in the cloud, edge servers, or end devices and (ii) the corresponding physical or software-form body. Every LM agent can actively interact with other LM agents, the virtual/real environment, and humans. For connected LM agents, there exist two typical communication modes: intra-agent communications for seamless data/knowledge synchronization between brain and physical body within an LM agent, and inter-agent communications for efficient coordination between LM agents. Table V summarizes the comparison of the two communication modes.
1) Intra-agent communications refer to the internal data/knowledge exchange within a single LM agent. This type of communication ensures that different components of the LM agent, including planning, action, memory, interaction, and security modules, work in harmony. For instance, an LM agent collects multimodal sensory data through its physical body, which then communicates the interpreted information to the LM-empowered brain. The planning module in the brain formulates a response or action plan, which is then executed by the action module. This seamless flow of information is critical for maintaining the LM agent’s functionality, coherence, and responsiveness in real-time and dynamic scenarios. These communications are semantic-aware, transmitting only contextually relevant data, such as changes in the environment.
For LM agents with physical embodiments, such as robots, ROS often serves as the foundational framework. ROS operates using a pub/sub mechanism, where components such as action module (e.g., sensors and actuators), planning module, and memory module either publish data streams or subscribe to them as needed. This structure enables flexible pub/sub interactions among various components, enabling LM agents to adapt quickly to dynamic environments. In summary, intra-agent communications will be semantic-aware and built on a pub/sub framework to align the LM agent’s internal processes.
2) Inter-agent communications involve information and knowledge exchange between multiple LM agents. It enables collaborative task allocation, resource sharing, and coordinated actions among agents to foster collective intelligence. For instance, in a smart city, various LM agents managing traffic lights, public transportation, and emergency services share real-time data to optimize urban mobility and safety. Effective inter-agent communications rely on standardized protocols to ensure compatibility and interoperability, facilitating efficient and synchronized operations across the LM agent networks. Two prerequisites should be met [25]: (1) a shared conceptual understanding of the subject matter being communicated, and (2) mutually intelligible encoding and decoding for accurately interpreting messages by both agents. Hence, agent-to-agent communications are also semantic-aware. Inter-agent communication protocols should at least specify the following key components:
  • Communication language, e.g., FIPA agent communications language (ACL) [87] and knowledge query and manipulation language (KQML) [88], defines the vocabulary and expressions used by agents to convey intentions.
  • Protocol syntax specifies the structure and format of messages, detailing essential fields like sender, receiver, content, and conversation ID.
  • Interaction strategies govern the sequences of communication, defining roles, turn-taking rules, and strategies for engaging in tasks. For instance, negotiation strategies dictate when agents can propose, counter, or conclude offers.
The design of inter-agent communication protocols faces the versatility-efficiency-portability trilemma: [25]:
  • Versatility: support diverse messages and applications to exchange information and coordinate actions.
  • Portability: operate across heterogeneous underlying platforms and support seamless integration of new agents and functionalities without significant redesign.
  • Efficiency: optimize resource usage and minimize communication and computational overheads.
For instance, traditional static APIs are highly efficient and portable but lack versatility, while natural language communication offers high versatility and portability but often sacrifices efficiency. LM agents bridge this gap by understanding, manipulating, and responding to other agents through natural language while integrating external tools, writing code, and invoking APIs. LM agents excel at following instructions, including implementing routines through code. Furthermore, they autonomously negotiate protocols and reach consensus on strategies and behaviors in complex scenarios. Marro et al. propose the Agora protocol [25] to addresses the communication trilemma by employing standardized routines for frequent communications, natural language for infrequent communications (e.g., negotiation and error handling), and LLM-written routines for scenarios that fall between them.
In addition to Agora, other agent communication protocols have been developed, e.g., Agent Protocol (adopted by AutoGPT), SmythOS, and Azure AI Agent Service. Selecting or designing an appropriate agent communication protocol requires careful consideration of task scenarios (e.g., wired backhaul or radio access networks) and task-specific requirements (e.g., robustness, latency, and cost constraints). Additionally, adapting these protocols to diverse systems and agent platforms presents significant challenges, highlighting the need for flexible and interoperable solutions.
Figure 12: Illustration of the cloud-edge-end cooperative network architecture of LM agents.
III-D Cloud-Edge-End Cooperative Framework for LM Agent Networking
III-D1 Illustrating Example in Autonomous Driving
Autonomous vehicles generate vast amounts of real-time data from sensors such as cameras and LiDAR, where LMs aid in making instant decisions on navigation, obstacle avoidance, and route optimization. For instance, LLMs can help determine the safest and most efficient route at intersections. However, autonomous driving demands low-latency responses, with 3GPP specifying end-to-end latency as low as 10 ms [89], which makes cloud-based LM processing unsuitable due to high latency. Furthermore, centralizing sensor data (up to 4TB/day [46]) in the cloud can overwhelm networks and compromise privacy, as sensitive information (e.g., location) may be exposed.
The cloud-edge-end collaborative framework addresses these issues by offloading LM processing to the edge, reducing bandwidth usage and ensuring faster response times. Edge or on-vehicle LM processing further protects privacy by avoiding data uploads to the cloud. This collaboration enhances autonomous driving through reduced latency, lower bandwidth costs, and improved privacy.
III-D2 Cloud-Edge-End Cooperative Framework Design
As depicted in Fig. 12, a cloud-edge-end cooperative framework is essential to resolve the above key challenges of LM agent networks. Generally, it includes three functional layers: cloud, edge, and end layers.
  • The cloud serves as the centralized intelligence hub, hosting full-scale LMs (e.g., exceeding 100B parameters) with extensive storage and computational resources. It performs computationally intensive tasks, such as large-scale model training, global decision-making, and cross-layer resource orchestration [90]. The cloud also maintains a global knowledge repository, which supports dynamic updates and synchronization across the network.
  • Edge nodes provide context-aware, intermediate intelligence by hosting moderately sized LMs (e.g., 10B–50B parameters) [30]. They bridge the gap between the cloud and end devices, performing task-oriented LM fine-tuning, proximal data aggregation, and contextual inference using localized datasets. They enable low-latency responses for region-specific tasks and support adaptive communications with end devices.
  • End devices (e.g., mobile agents and IoT devices) execute lightweight, real-time tasks using tiny local LMs (e.g., 0–10B parameters). They provide contextual inputs, collect real-time sensory data, and act as the first responders in scenarios requiring immediate action. For instance, Google has launched the Gemini Nano [^11] https://store.google.com/ideas/gemini-ai-assistant/ on Pixel 8 Pro smartphones with 1.8 billion and 3.25 billion parameters, which support relatively basic functions such as grammar checking and text summarization. End devices also prioritize privacy preservation by processing sensitive data locally before sharing insights with edge or cloud layers.
For instance, Xu et al. [30] propose an edge computing framework tailored for LLM agents. In their framework, mobile agents with local models (0-10B parameters) handle real-time tasks, and edge agents with larger models (over 10B parameters) process broader contextual data to support complex decision-making. They also study a real use case of vehicle accidents, where mobile agents create localized accident scene descriptions, which are then enhanced by edge agents to generate comprehensive accident reports and actionable plans.
LM lightweighting techniques, e.g., knowledge distillation [26], quantization [27], pruning [28], and hardware acceleration [29], can be employed to deploy compact LM variants tailored to edge/end computational capacities. For example, VPGTrans [26] addresses the high computational costs of training visual prompt generators for multimodal LLMs by significantly reducing GPU hours and training data needed via transferring an existing visual prompt generator from one multimodal LLM to another. LLM-Pruner [28] compresses LLMs without compromising their multi-task solving and language generation abilities, using a task-agnostic structural pruning method that significantly reduces the need for extensive retraining data and computational resources. Hardware accelerators, such as GPUs and TPUs, can enhance computational efficiency and provide robust support for cloud training and edge deployment. Agile-Quant [29] enhances the efficiency of LLMs on edge devices through activation-guided quantization and hardware acceleration by utilizing a SIMD-based 4-bit multiplier and efficient TRIP matrix multiplication, achieving up to 2.55x speedup while maintaining high task performance.
III-D3 Advantages
By assigning specialized roles to cloud, edge, and end devices, the cloud-edge-end architecture enhances flexibility to accommodate varying task complexity and resource availability. For instance, computation-intensive LM tasks can be offloaded to cloud servers, less demanding LM tasks can be handled by edge servers to reduce latency and bandwidth usage, while end devices can offload latency-sensitive LM operations to edge servers. Additionally, hierarchical collaborative learning is essential for LM agents to acquire global knowledge across large geographical areas, with LM updates of edge/end agents synchronized centrally in the cloud. This hierarchical collaboration across cloud, edge, and end devices offers several key advantages:
  • Scalability: By distributing computational workloads across layers, the architecture supports large-scale networks with heterogeneous agents.
  • Efficiency: The cloud handles high-complexity tasks, while edge and end devices manage localized, latency-sensitive operations, optimizing overall system performance.
  • Adaptability: Dynamic coordination across layers ensures that the system can respond flexibly to changes in spatiotemporal dynamics, such as agent mobility and evolving tasks.
  • Security and Privacy: Localized processing at the end layer reduces the exposure of sensitive data, while the cloud and edge nodes provides robust security mechanisms for cross-layer interactions.
III-E Interaction Strategies of Connected LM Agents
The multi-agent interactions can be categorized into distributed cooperation, distributed competition, and partial cooperation [82, 63], each of which involves different strategies to jointly optimize the collective or individual outcomes.
III-E1 Distributed Cooperation
For connected LM agents, cooperation is a fundamental interaction mode where distributed LM agents work together to achieve common goals, share resources, and optimize collective outcomes. It includes three aspects: a) distributed cooperative LM fine-tuning, b) distributed cooperative LM inference, and c) distributed cooperative LM caching.
Figure 13: Illustration of (a) split learning for collaborative LM Fine-tuning and (b) split inference for collaborative LM inference in edge environments.
a) Distributed Cooperative LM Fine-tuning for LM Agents. LMs, typically designed for general tasks, require fine-tuning with vast domain-specific data on edges to provide agent services in specialized fields. However, due to their massive size, fine-tuning LMs demands substantial computational power and energy, as shown in Table IV. To fine-tune LMs on heterogeneous, resource-constrained edge environments while ensuring privacy protection, we introduce two paradigms: collaborative split learning (SL) and federated edge learning, as below.
  • Cooperative SL. SL distributes partitioned LMs on edge devices and servers. In SL, as shown in Fig. 13(a), edge devices upload data features to edge servers for co-inference, reducing local computation while retaining major processing tasks at the edge. By placing only a submodel on edge devices, SL minimizes their computational burden and can integrate with parameter-efficient fine-tuning (PEFT) techniques [91] to further reduce workload. However, deploying SL for LM agents in edge-end environments introduces several challenges. First, communication costs for transmitting high-dimensional data can be significant. For example, partitioning large models generates smashed data that should be uploaded, resulting in substantial overhead. In the case of GPT-3 Medium and 100 samples (1024 tokens each), this can reach approximately 400 MB per training round [92]. Second, privacy concerns arise due to target token leakage. In common edge SL configurations, the input module is placed on the edge and the output module on the server, requiring edge devices to upload target tokens for training [93]. This can expose sensitive user data and compromise privacy. U-shaped SL [94] retains the initial and final neural layers or transformer blocks on edge devices, placing only intermediate layers on edge servers, thereby safeguarding label privacy to a large extent.
  • Federated edge learning. In cloud-edge collaborative LM training, LMs are initially pre-trained in the cloud, utilizing its extensive computational resources and large-scale datasets. Federated fine-tuning technology is then performed on edge devices using domain-specific data from various edge servers, enabling LM agent services tailored to specific requirements of diverse localized downstream tasks and optimizing resource utilization across the network. However, due to the massive size of LMs (often containing billions of parameters), uploading model updates from numerous edge servers for aggregation during the federated fine-tuning process imposes significant communication overheads and operational costs. Additionally, periodic LM local fine-tuning at the edge consumes substantial computational resources, which often exceed the capacity of resource-constrained edge nodes. The offsite-tuning transfer learning technique, in conjunction with PEFT techniques [91], such as Prefix-tuning and LoRA, enable edge devices to update and transmit only a subset of model parameters, thereby significantly reducing both communication and computational overheads [85]. Specifically, as shown in Fig. 14, the cloud employs offsite-tuning transfer learning technique to generate a lightweight proxy model that encodes task-specific knowledge with a minimal number of parameters. Subsequently, edge nodes fine-tune the PEFT module of this proxy model to further reduce the communication and computational costs.
Figure 14: Illustration of federated edge learning in cloud-edge environments for collaborative LM Fine-tuning.
Figure 15: Illustration of MoE-oriented collaborative LM inference for LM agents in cloud-edge-end environments.
b) Distributed Cooperative LM Inference for LM Agents. On-server LM inference offloads computation to cloud servers but compromises user privacy by requiring raw data uploads. On-device inference, on the other hand, preserves privacy but imposes heavy computational burden on edge devices. To resolve this dilemma in resource-constrained environments, mixture-of-experts (MoE) and split inference are two key approaches. Additionally, LM agents can leverage role-playing for cooperative inference in general applications.
  • Cooperative LM inference via MoE. MoE enhances large models (LMs) by partitioning them into specialized expert networks, each tailored to process specific input types [31]. A gating network dynamically assigns tokens to the most appropriate expert(s), enabling sparse activation that reduces unnecessary computations and improves efficiency. MoE has been shown to boost performance by 45% while consuming only one-third of the resources required by dense models [95]. In cloud-edge architectures, as shown in Fig. 15, cloud servers store the expert models, and edge servers and devices download only the relevant experts for local inference. This method optimizes resource utilization, accelerates processing, and makes the deployment of large models viable in resource-constrained environments. For instance, EdgeMoE [96] stores non-expert weights in memory while placing larger expert weights on disk, activating only the necessary experts. This approach reduces memory consumption by 2.6 to 3.2 times compared to full model loading. However, a series of challenges remain in MoE edge deployment, as follows. 1) MoE models are primarily designed for high-performance cloud data centers equipped with powerful GPUs and ample network bandwidth. However, when deployed on edge devices, MoE models must accommodate the wide variety of hardware capabilities, ranging from basic IoT devices to advanced edge servers. This heterogeneity requires flexibility in MoE design to ensure high performance while adapting to the varying resource constraints of edge environments. 2) Edge environments are highly dynamic, which complicates inference tasks. When local expert networks are unavailable, they must be dynamically downloaded, a process influenced by network speed and device capacity. Additionally, selecting the appropriate expert network adds complexity to model verification and management. 3) As the expert networks required for inference tasks may also expose users’ environmental context and behavioral patterns, it is essential to incorporate privacy protection measures when designing the inference engine.
  • Cooperative split inference. Split inference, as outlined in the 3GPP 5G specifications [89], partitions an LM between edge devices and edge/cloud servers, offloading major part of computation to edge/cloud servers for mitigated device-side workload. It offers a middle ground between on-server and on-device inference. As shown in Fig. 13(b), edge devices perform initial inference using the user-side submodel and upload intermediate features to the server for further processing, thereby optimizing resource usage while preserving privacy by keeping sensitive raw data on the device. According to [46], split inference for LLMs can be implemented in two ways: 1) In encoder-decoder LMs such as BART [97], the encoder is placed on the device, while the more computationally demanding decoder is on the server. 2) In decoder-only LMs such as GPT-series, lightweight components (e.g., embeddings, early Transformer blocks) stay on the device, while heavier components are offloaded to the server. Additionally, critical feed-forward network (FFN) parameters can be stored on the device to save memory. However, split inference faces challenges such as high communication costs due to large data uploads, latency from reliance on both edge devices and servers, and privacy concerns from potential leakage of intermediate outputs. Early exit strategies help lower computational latency on edge devices and servers. Transformer outputs can further be compressed via quantization, pruning, and merging [27]. Additionally, progressive split inference [98] can eliminate unnecessary transmission of intermediate token representations while maintaining desired inference accuracy.
  • Role-playing. Role-playing [99, 67] is a typical method for facilitating cooperation among LM agents. 1) Instruction-following cooperation. Li et al. [100] propose a cooperative agent framework that employs role-playing with inception prompting to guide agents autonomously toward task completion. The system begins with human-provided ideas and role assignments, refined by a task-specifier agent. An AI user and an AI assistant then collaborate through multi-turn dialogues, with the AI user instructing and the assistant responding until the task is completed. 2) Instruction-free cooperation. Fig. 16 shows an example of role-playing based novel writing using collaborative agents [101]. ChatDev [67] is a virtual software company operated by “software agents” with diverse roles, e.g., chief officers, programmers, and test engineers. Following the waterfall model, it divides development into design, coding, testing, and documentation, breaking each phase into atomic subtasks managed by a chat chain. To create cooperative agents with adaptive behaviors, ProAgent [84] assigns each agent three roles: planner, verificator, and controller. The planner formulates high-level skills, incorporating belief revision and knowledge management. The verificator assesses the feasibility of skills, identifies issues if a skill fails, and initiates re-planning when needed. If the skill is feasible, the controller decomposes it into executable actions. Park et al. [99] create a community of 25 generative agents in a sandbox world named Smallville, where agents are represented by sprite avatars. These agents perform daily tasks, form opinions, and interact, mimicking human-like behaviors.
Figure 16: An example of role-playing for novel writing using collaborative agents [101]. In the draft stage, the Planner creates an initial agent team and action plan, refining it through feedback from the Agent and Plan Observers. The Action Observer assigns tasks, monitors execution, and adjusts the plan based on performance. Agents collaboratively refine and execute tasks, adapting as needed.
c) Distributed Cooperative LM Caching for LM Agents. Edge model caching can significantly reduce model download latency by pre-distributing LMs to wireless edge servers, allowing end users to download them directly. This strategy provides users with quick access to LMs within their quality-of-service (QoS) requirements, bypassing the need to retrieve them from remote cloud servers, which would incur high latency.
  • Edge LM Caching. Due to the limited storage capacity, edge servers need to optimize the cache hit ratio by placing popular LMs on edge servers. For instance, TrimCaching [102] optimizes LLM model placement by leveraging LLM parameter sharing to maximize cache efficiency under storage and latency constraints. By storing a single copy of shared parameter blocks on each edge server, it enhances storage utilization. Practical edge LM caching deployment also needs to consider user mobility patterns and inter-edge collaboration and competition strategies under different LM partitioning methods (e.g., MoE and SL).
  • Edge LM Replacement. As user demands evolve, previously cached LMs may no longer match current request patterns. In such cases, edge servers should replace older models with more relevant ones. This process incurs high communication costs, as transferring large-scale models can impose substantial overhead on mobile backhaul networks. Recency-based and frequency-based are two classic strategies, replacing the least recently used (LRU) or least frequently used (LFU) items with updated ones. Popularity-based methods, such as SwapMoE [103], prioritize caching highly activated expert networks, with importance determined by activation frequency during inference tasks. However, these approaches often neglect shared parameter blocks across LLMs and the potential for cooperative caching among edge nodes. Distributed reinforcement learning can be leveraged to make dynamic replacement decisions without requiring complete information from other edge nodes.
  • Edge LM Routing. To further minimize service latency, it is essential to optimize edge LM routing strategies that direct users to the most suitable edge server based on network conditions, queue length, and model availability. Additionally, due to the mobility of agents, it requires seamless model migration across edge servers to maintain optimal service delivery. Further research efforts are required for intelligently managing user requests and distributing LM workloads across edge nodes to enhance performance, minimize latency, and maximize resource utilization.
  • Edge Caching for RAG. LM agents’ performance is often tightly linked to the quality and relevance of the knowledge they possess. Therefore, selecting which data to cache for LM updates via RAG is essential to ensure that LM agents provide reliable, up-to-date services. Efficient knowledge cache management not only improves response times but also supports the continuous learning and adaptation of LM agents in dynamic environments. Further research is needed to tackle the joint LM caching and knowledge caching problem within the context of RAG under practical constraints of edge environments.
Figure 17: Illustration of (a) one-by-one debate, (b) simultaneous-talk debate, and (c) simultaneous-talk-with-summarizer debate [104].
III-E2 Distributed Competition
Competition involves LM agents pursuing their individual objectives, often at the expense of others. This interaction mode is characterized by non-cooperative strategies where agents aim to maximize their own benefits. Multi-agent debate (MAD) and non-cooperative game have been widely adopted to model the competitive behaviors between autonomous agents.
  • MAD. The cognitive behavior of LLMs, such as self-reflection, has proven effective in solving NLP tasks but can lead to thought degeneration due to biases, rigidity, and limited feedback. LM agents in MAD [104] involve engaging in structured arguments or debates to defend their positions and challenge the strategies of others. MAD explores interactions among LLMs, where agents engage in a dynamic tit-for-tat, allowing them to correct each other’s biases, overcome resistance to change, and provide mutual feedback. In MAD, diverse role prompts (i.e., personas) are crucial and there are generally three communication strategies: (a) one-by-one debate, (b) simultaneous-talk, and (c) simultaneous-talk-with-summarizer, as illustrated in Fig. 17. In one-by-one debate strategy, debater agents respond sequentially in a fixed order, generating replies based on the ongoing discussion. Each agent’s chat history is updated with previous responses to maintain context. In simultaneous-talk debates, debater agents respond asynchronously within each round, mitigating the influence of speaking order on the debate’s flow. The simultaneous-talk-with-summarizer debate strategy enhances simultaneous-talk by incorporating a summarizer LLM. After each round, the summarizer condenses the discussion into a cohesive summary, which replaces the chat histories of all agents, ensuring shared context.
  • Non-cooperative Game. In non-cooperative games, LM agents engage in strategic decision-making where each LM agent’s goal is to seek the Nash equilibrium. Various non-cooperative games can be employed such as Nash bargaining, auctions, matching games. For instance, Xu et al. [105] design a new cached model-as-a-resource paradigm and introduce the concept of age-of-thought to efficiently deploy LLM agents. They also devise a deep Q-network-based auction mechanism to incentivize network operators.
III-E3 Partial Cooperation
Partial cooperation occurs when LM agents collaborate to a limited extent, often driven by overlapping but not fully aligned interests [63]. In such scenarios, agents might share certain resources or information while retaining autonomy over other aspects. This interaction mode balances the benefits of cooperation with the need for individual agency and competitive advantage. Partial cooperation can be strategically advantageous in environments where complete cooperation is impractical or undesirable due to conflicting goals or resource constraints. For instance, DyLAN [106] is a multi-layered LLM agent network for collaborative task-solving with teammates, such as reasoning and code generation. DyLAN facilitates multi-round interactions in a dynamic setup with an LLM-empowered ranker to deactivate low-performing agents, an early-stopping mechanism based on Byzantine consensus to efficiently reach agreement, and an automatic optimizer to select the best agents based on agent importance scores.
Hierarchical game and coalition formation theory can be employed to model partially cooperative interactions among LM agents. Hierarchical game [107] structures the interactions of LM agents across different game stages, where they may cooperate at one level and compete at another. Through the coalition formation theory [83], LM agents can form the optimal stable coalition structure under different scenarios and environments, where the optimal stable coalition structure is featured with intra-coalition cooperation and inter-coalition competition.
III-E4 Lessons Learned
In distributed collaboration, SL and MoE focus on model segmentation, while FL and role-playing emphasize model integration. Within the overall cloud-edge collaboration for LM agent systems, key challenges emerge in the joint consideration of LM segmentation (e.g., SL and MoE), LM placement, LM replacement, LM routing, and knowledge caching, alongside the inherent heterogeneity and dynamics of edge environments. Effective deployment and operation of LM agents across cloud and edge infrastructures require a delicate balance between LM segmentation and LM integration strategies, with an emphasis on optimizing resource allocation and minimizing service latency. Additionally, solutions should be adaptable to the constantly changing network conditions and hardware limitations inherent to edge devices. Collaborative models such as SL and MoE can significantly reduce computational burdens on edge devices, while federated strategies and role-playing can enhance cooperation among distributed agents. Future research should explore the seamless integration of these techniques, taking into account both static and dynamic aspects of edge networks, to provide more efficient and scalable solutions for large-scale applications.
Figure 18: Illustration of the cooperation modes of connected LM agents including data cooperation, computation cooperation, and knowledge cooperation.
III-F Cooperation Paradigms of Connected LM Agents
As illustrated in Fig. 18, the detailed cooperation paradigms of LM agents under cloud-edge-end architecture involves three perspectives: data cooperation, computation cooperation, and knowledge cooperation.
  • Data Cooperation: Within a common task, LM agents continuously exchange and fuse their individual data (e.g., task-oriented sensory data) to ensure a comprehensive and up-to-date understanding of their task environment, thereby enhancing collective intelligence and enable coordinated actions.
  • Computation Cooperation: LM agents perform coordinated reasoning, such as service cascades, by distributing computational tasks optimally across agents, leveraging collective processing power to handle complex computations more efficiently.
  • Knowledge Cooperation: LM agents share domain-specific knowledge and experiences (e.g., in the format of KGs) to collectively improve problem-solving capabilities for better-informed actions and decisions, via knowledge synchronization, fusion, and retrieval and distributed learning algorithms.
Figure 19: Illustration of (a) horizontal collaboration and (b) vertical collaboration paradigms for multiple LM agents.
1) Data Cooperation for LM Agents: The data cooperation among LM agents involves the modality perspective and the spatio-temporal perspective.
a) Multimodal Data Cooperation emphasizes the fusion of data from various modalities, such as text, images, audio, and video, to provide a comprehensive understanding of the environment. This cooperation allows LM agents to process and interpret information from multiple sources, leading to more accurate and robust decision-making. (i) By combining data from different modalities, it helps to create a unified representation that takes advantage of the strengths of each type of data. For instance, Gross et al. [108] discuss the use of multimodal data to model communication in artificial social agents, emphasizing the importance of verbal and nonverbal cues for natural human-robot interaction. (ii) By enabling LM agents to retrieve relevant information across different modalities, it enhances their ability to respond to complex queries and scenarios. For instance, by considering the intra-modality similarities in multi-modal video representations, Zolfaghari et al. [109] introduce the contrastive loss in contrastive learning process for enhanced cross-modal embedding. The effectiveness is validated using LSMDC and Youcook2 datasets for video-text retrieval task and video captioning task.
b) Spatio-temporal Data Cooperation involves the integration and synchronization of spatial and temporal data across various modalities and sources, enabling LM agents to achieve a comprehensive and dynamic understanding of the environment over time. This cooperation ensures that LM agents can effectively analyze patterns, predict future states, and make informed decisions in real-time, based on both spatial distribution and temporal evolution of data. Yang et al. [86] introduce SCOPE, a collaborative perception mechanism that enhances spatio-temporal awareness among on-road agents through end-to-end aggregation. SCOPE excels by leveraging temporal semantic cues, integrating spatial information from diverse agents, and adaptively fusing multi-source representations to improve accuracy and robustness. However, [86] mainly works for small-scale scenarios. By capturing both spatial and temporal heterogeneity of citywide traffc, Ji et al. [110] propose a novel spatio-temporal self-supervised learning framework for traffic prediction that improves representation of traffic patterns. This framework uses an integrated module combining temporal and spatial convolutions and employs adaptive augmentation of traffic graph data, supported by two auxiliary self-supervised learning tasks to improve prediction accuracy.
2) Computation Cooperation for LM Agents: It can be classified into three modes: horizontal cooperation, vertical cooperation, and hybrid cooperation.
a) Horizontal collaboration refers to decompose complex tasks into manageable sub-tasks, with multiple LM agents independently completing their assigned tasks in parallel. The results are then summarized and integrated to generate the final outcome. This approach enables the collaborative system to scale horizontally by adding more LM agents, allowing for more complex and dynamic tasks to be handled. The parallel processing and multi-angle perspectives contribute to increased robustness and reduced errors and biases that may arise from relying on a single agent. As shown in Fig. 19(a), each LM agent independently analyzes different aspects of stock performance such as trend, news, and industry to assess its investment potential. Horizontal collaboration has been successfully implemented in various LM agent systems. For instance, ProAgent [84] employs a framework where LM agents collaborate as teammates, analyzing each other’s intentions and updating their beliefs based on observed behaviors of their peers. This enhances collective decision-making by allowing agents to adapt dynamically in real time. Similarly, DyLAN [106] assembles a team of strategic agents that communicate through task-specific queries, allowing multiple rounds of interaction to enhance both efficiency and overall performance. However, the independent analysis of each LM agent may lead to inconsistent outputs, complicating the aggregation of the final conclusions. Therefore, effective mechanisms are needed to address disagreements and ensure that the agents’ contributions are complementary, which may increase the complexity of the coordination process.
b) Vertical collaboration involves decomposing complex tasks into multiple stages, with different LM agents handling each stage sequentially. After completing their respective tasks, each agent passes the result to the next agent, until the entire task is successfully completed. As illustrated in Fig. 19(b), in medical image analysis, the first agent might extract basic visual features, the second agent could analyze these features in greater depth, and the final agent produces a diagnosis. This sequential approach enables a step-by-step refinement, allowing complex problems to be broken down and tackled more efficiently by leveraging specialized agents at each stage. For instance, Jiang et al. [24] present CommLLM, a multi-agent system for natural language-based communication tasks. It comprises three components: 1) multi-agent data retrieval, which uses condensate and inference agents to refine 6G communication knowledge; 2) multi-agent collaborative planning, employing various planning agents to generate solutions from multiple viewpoints; and 3) multi-agent reflection, which evaluates solutions and suggests improvements through reflexion and refinement agents. However, a limitation of vertical collaboration is that higher-level LM agents rely on the accuracy of lower-level outputs, making the process vulnerable to error propagation. Mistakes made early in the sequence can compromise the final outcome, highlighting the need for high precision at each stage to maintain the overall system’s reliability.
c) Hybrid collaboration. In practical LM agent environments, real-world applications often require a combination of horizontal and vertical collaboration, resulting in hybrid collaboration, as illustrated in Fig. 16. For instance, when addressing highly complex tasks, the problem is first broken down into manageable sub-tasks, each assigned to specialized LM agents. Horizontal collaboration allows agents to perform parallel evaluations or validations, while vertical collaboration ensures that the task is refined through sequential processing stages. This computational collaboration blends both paradigms, coordinating parallel assessments across agents while sequentially integrating their outputs through defined stages, thus optimizing task execution and enhancing overall system performance. For instance, Chen et al. [111] propose a dynamic multi-agent collaboration framework that organizes agents using directed acyclic graphs to facilitate interactive reasoning. Their framework demonstrates superior performance across various network topologies and enables collaboration among thousands of agents. A key finding in [111] is the discovery of the collaborative scaling law, where solution quality improves in a logistic growth pattern as more agents are added, with collaborative emergence occurring faster than neural emergence.
3) Knowledge Cooperation for LM Agents: Knowledge can be broadly categorized into explicit knowledge and implicit knowledge. Explicit knowledge refers to structured, codified, and easily accessible information that can be articulated, documented, and shared, such as external databases and KGs. Implicit knowledge is embedded within the model and arises through the optimization of its internal parameters, such as weights and biases, during training or fine-tuning [112]. The knowledge cooperation generally involves three consecutive aspects: knowledge synchronization, knowledge fusion, and knowledge retrieval.
a) Knowledge sharing. It includes the sharing and updating of knowledge between multiple LM agents to ensure consistency in decision making.
① Knowledge transfer is a common method of knowledge sharing that transfers the implicit knowledge, in the form of learned parameters, from one LM agent to another. Kang et al. [113] propose an online distillation method that enhances LLMs by retrieving relevant knowledge from external knowledge bases, to generate high-quality reasoning processes. The knowledge (e.g., reasoning results) of small language models can be leveraged to improve the performance of LLMs in knowledge-intensive tasks. However, as the scale of LLMs grows, LLM training via such knowledge transfer becomes complex and computationally intensive. To address this issue, Zhong et al. [114] propose a parametric knowledge transfer method that extracts and aligns knowledge parameters using sensitivity techniques and uses the LoRA module as an intermediary mechanism to inject this knowledge into smaller models, transferring the implicit knowledge of smaller models.
② Knowledge update. Knowledge alignment is a prerequisite to update knowledge among LM agents. Zhang et al. [115] propose a fully automatic KG alignment method, using LLMs to identify and align entities in different KGs, aiming to solve the heterogeneity problem between different KGs and integrate multi-source data. (i) For explicit knowledge updates, Tandon et al. [116] pair LMs with a growing memory to train a correction model, where users identify output errors and provide general feedbacks on how to correct them. (ii) For implicit knowledge updates, continual learning ensures that AI model can continuously learn while receiving new tasks and new data without forgetting previously learned knowledge. Qin et al. [117] propose ELLE, which flexibly extends the breadth and depth of existing PLMs, allowing the model to continuously grow as new data flows in.
b) Knowledge fusion. It contains the knowledge fusion and completion phases between multiple LM agents.
① Knowledge integration, which integrates and optimizes knowledge from different LM agents to form a robust and comprehensive public knowledge base for a common task. Jiang et al. [118] propose an ensemble framework named LLM-blender that directly aggregates the outputs of different models to improve prediction performance and robustness. However, it requires maintaining multiple trained LLMs and executing each LLM during inference, which can be impractical for LM agents. To address this issue, Wan et al. [79] propose an implicit knowledge fusion framework that evaluates the predictive probability distributions of multiple LLMs and uses the distribution for continuous training the target model.
② Knowledge completion, which infers missing facts in a given KG. Compared with traditional KG completion methods, LLMs hold the potential to enhance KG completion performance via encoding text or generating facts. Shen et al. [119] use LLMs as encoders, primarily capturing the semantic information of KG triples through the model’s forward pass, and then reconstructing the KG’s structure by calculating a loss function, thereby better integrating semantic and structural information. Apart from KG completion, LLMs can be employed to verify KGs. Han et al. [120] propose a prompt framework for iterative KG verification, using small LLMs to correct errors in KG generated by LLMs such as ChatGPT.
TABLE VI: A Summary of Knowledge Cooperation Approaches for LM Agents
Stage Type Technology Ref.
Online knowledge distillation LMs generate high-quality reasoning processes for models to learn. [113]
Offline knowledge distillation Extract and align parameters of LLMs to smaller models. [114]
Knowledge Sharing Explicit knowledge update Adjust LM via knowledge bases and feedbacks. [116]
Implicit knowledge update Selective fine-tuning of model’s internal parameter knowledge. [117]
Implicit knowledge fusion Train target model via predictive probability distributions of LMs. [79]
Model output fusion Directly aggregate the outputs of different models. [118]
Knowledge Fusion KG completion Encode text to better use textual & semantic infomation. [119]
KG verification Using a small LM to validate and correct the LM output. [120]
Knowledge Retrieval Static knowledge base Retrieve knowledge from static sources for accurate reasoning. [58, 59]
Dynamic knowledge base
Use real-time updated knowledge sources for accurate reasoning.
[80, 78, 77]
c) Knowledge retrieval. LM agents not only rely on the learned knowledge during pre-training but also can dynamically access and query external knowledge bases (e.g., databases, the Internet, and KGs) to obtain the latest information to help reasoning. RAG technology combines information retrieval and generation models by first retrieving relevant contents and then generating answers based on these contents. According to the data type, it can be divided into two categories.
  • RAG based on static knowledge sources, such as Wikipedia and documents. Lewis et al. [58] demonstrate how RAG generates accurate and contextually relevant responses by retrieving relevant knowledge chunks from static sources. For retrieval enhancement, Trivedi et al. [59] propose a new multi-step QA method named IRCoT to interleave retrieval with steps in CoT, which first uses CoT to guide retrieval and then uses retrieval results to improve CoT.
  • RAG based on dynamic knowledge sources, such as news APIs. It contains three lines: exploring new knowledge, retrieving past knowledge, and self-guided retrieval. (i) For new knowledge exploration, Dai et al. [80] use RAG technology for improved safety and reliability of autonomous driving systems by utilizing real-time updated data sources, including in-vehicle sensor data, traffic information, and other driving-related dynamic data. It demonstrates RAG’s potential in handling complex environments and responding to unexpected situations. (ii) For past knowledge retrieval, Kuroki et al. [78] develop a novel vector database named coordination skill database to efficiently retrieve past memories in multi-agent scenarios to adapt to new cooperation missions. (iii) For self-guided retrieval, Wang et al. [77] propose a method called self-knowledge-guided retrieval (SKGR), which harnesses both internal and external knowledge for enhanced retrieval by allowing LLMs to adaptively call external resources when handling new problems.
III-G Summary and Lessons Learned
LM agent networks bridge the human, physical, and cyber worlds, with an architecture that integrates cloud, edge, and end devices. Powered by advanced technologies including foundation models, knowledge-related technologies, interaction technologies, multi-agent collaboration, and digital twin technologies, LM agent networks enable seamless communication, sophisticated reasoning, and scalable task execution. By leveraging hierarchical frameworks, semantic-aware communication, information-centric routing, and spatiotemporal awareness, LM agent networks optimize resource allocation, enhance real-time decision-making, and improve both intra- and inter-agent communication.
  • Intra-agent communication allows different components of a single LM agent, such as planning, action, memory, and security modules, to function cohesively. Frameworks such as the pub/sub mechanism facilitate efficient data exchange based on relevance. Inter-agent communication enables multiple LM agents to collaborate, share resources, and coordinate actions for collective intelligence. Protocols such as FIPA ACL and KQML define the vocabulary, message structure, and interaction strategies essential for effective communication between agents. Communication protocols should align with network environments (e.g., wireless or wired), latency requirements, robustness, and security needs. Adapting these protocols to various systems and agent platforms presents a significant challenge, highlighting the need for flexible, interoperable solutions.
  • The cloud-edge-end architecture allows for scalable and efficient task allocation, addressing the varying computational capabilities and real-time requirements of different devices.
  • Horizontal, vertical, and hybrid collaborations provide complementary solutions to complex problems, combining distributed and sequential task processing to optimize overall system performance. This includes cooperative LM fine-tuning, where agents collaboratively adjust LMs using domain-specific data; cooperative LM inference, where inference tasks are distributed across agents to speed up processing; and collaborative edge LM caching, where LM (or LM parts) are stored and shared across edge devices to reduce latency and improve access efficiency.
  • The synchronization of implicit and explicit knowledge is crucial for enhancing the adaptability and reliability of LM agents, enabling them to respond effectively in dynamic, evolving environments. Implicit knowledge, often derived from experience and context, allows agents to quickly adapt to unforeseen situations, while explicit knowledge, i.e., structured, defined information, ensures agents can make informed decisions based on clear rules or facts.
Figure 20: The taxonomy of security threats to LM agent networks.
IV Security Threats & Countermeasures to Large Model Agent Networks
In previous sections, we have explored the core concepts of LM agents and the importance of their networks for efficient collaboration and data/knowledge sharing. However, the three-tiered decentralized nature of LM agent networks, while offering scalability and flexibility, presents significant challenges related to secure and reliable collaboration throughout the life cycle of LM agent services. In this section, we present a comprehensive review of security threats related to LM agent networks and examine the state-of-the-art countermeasures to defend against them. Fig. 20 illustrates the taxonomy of security threats to LM agent networks.
Figure 21: Illustration of authentication threats in LM agent networks: (a) user-agent authentication, (b) agent-agent authentication threats, and (c) agent-TPSP (third-party service provider) authentication.
IV-A Authentication Threats in LM Agent Networks
Efficient and reliable authentication mechanisms is a prerequisite for safeguarding the functionality and integrity of LM agent systems. Based on the interacting entities, authentication threats to LM agents can be categorized into three types: user-agent authentication threats, agent-agent authentication threats, and agent-TPSP (third-party service provider) authentication threats, as illustrated in Fig. 21.
IV-A1 User-Agent Authentication Threats
LM agents generally store a wealth of private data for users and extensive sensitive corporate information for enterprises, making them prime targets for adversaries. Attackers can exploit identity forgery, credential theft, authentication bypassing, or privilege abuse to gain unauthorized access, leading to data breaches, system misuse, or service disruptions. Unlike traditional authentication scenarios, adversaries can leverage the advanced generative capabilities of LM agents to enhance their ability to bypass authentication. For instance, as illustrated in Fig. 21(a), an attacker could deploy an AIGC agent to forge a digital identity, imitating a legitimate user’s voice, appearance, and behavior using publicly available media from social platforms to evade multi-factor authentication (MFA). If successful, the attacker could exploit the vast personal data within LM agents to construct a digital replica of the victim or access critical corporate assets, resulting in greater losses than traditional authentication threats. Furthermore, as privatized LM agents become increasingly prevalent, interaction contexts are likely to contain even more personal and sensitive user information. This amplifies the impact of authentication threats, potentially leading to severe privacy violations and significant damage.
IV-A2 Agent-Agent Authentication Threats
As LM agents increasingly collaborate to tackle complex tasks through horizontal, vertical, or hybrid coordination, authentication between LM agents becomes critical to LM agent networks. A primary threat in such scenarios is agent spoofing attacks, where adversaries compromise or impersonate legitimate agents within the collaboration workflow. For instance, in horizontal or vertical collaboration scenarios, attackers may compromise LM agents, converting them into malicious entities, or impersonate legitimate agents to disrupt collaborative LM inference processes, thereby undermining the accuracy and reliability of the collaboration system and jeopardizing the task’s outcome, as illustrated in Fig. 21(b). Additionally, collusive LM agents can sabotage the consensus-building process required for task execution, introducing misinformation or bias that disrupts the overall functionality and trustworthiness of the collaborative system.
IV-A3 Agent-TPSP Authentication Threats
LM agents often rely on external tools, web APIs, and third-party service providers (TPSPs) to execute diverse tasks, making secure and reliable authentication between agents and TPSPs critical. This process involves mutual verification: (i) TPSPs should confirm the agent’s legitimacy and authorization, typically via a valid API key or token; (ii) agents should authenticate TPSPs to ensure they interact with trusted and legitimate service endpoints. Authentication threats in this context primarily target the communication channel between LM agents and TPSPs, exposing them to interception-based attacks such as man-in-the-middle (MITM) and session hijacking. For instance, an adversary could intercept agent-TPSP communications, altering responses or injecting erroneous data, leading to unexpected outputs for users, as illustrated in Fig. 21(c). Additionally, improper API key or token management can exacerbate risks and result in unauthorized service misuse.
IV-A4 Countermeasures to Authentication Threats in LM Agent Networks
While existing technologies such as end-to-end encryption, certificates, MFA, and key management are effective in mitigating authentication threats in LM agent services, their effectiveness diminishes in the face of the growing complexity of LM agent collaboration and the evolving sophistication of attacks driven by LMs. Moreover, the rapid development of LM agents has created a dual-edged dynamic, where both attacks and defenses are increasingly automated, escalating the risks and complexities of maintaining security. As a result, it is crucial to develop adaptive and collaborative resilient authentication mechanisms to ensure the security and reliability of LM agent systems in complex, dynamic environments.
  • Risk-Based Adaptive Authentication: This approach leverages the contextual reasoning capabilities of LM agents to dynamically assess the risk of user interactions. By analyzing user behavior, intent, and contextual anomalies, the authentication risk of the entity being authenticated is evaluated in real-time. Based on the risk level, the authenticator (such as the user, LM agent, or TPSP) adapts the authentication requirements, which may include triggering additional verification steps or involving third-party trusted institutions. This adaptive and flexible mechanism enhances the responsiveness of LM agents to varying levels of risk.
  • Multi-Agent Cooperative Authentication: This approach is designed for scenarios requiring long-term collaboration between LM agents. For instance, when multiple LM agents are working together to complete a task, each agent independently verifies the outputs or requests of the target agent for potential anomalies. The output is integrated into the collaborative workflow only after a sufficient number of agents confirm its reliability, based on a predefined threshold. This cooperative verification process reduces the risks of single-point failures, strengthens the robustness of authentication in dynamic agent networks, and mitigates the likelihood of adversarial manipulation.
Figure 22: Illustration of DoS atttacks in LM agent networks: (a) sponge examples and (b) intelligent botnet.
IV-B Denial-of-Service (DoS) Attacks in LM Agent Networks
The fine-tuning and inference processes of LM agents require significant computational resources. DoS attacks, however, can drastically increase resource consumption, overwhelming LM agents and disrupting service availability. Based on the attack method, DoS attacks in LM agent networks can be classified into two types: data-oriented DoS attacks (e.g., sponge examples) and flooding-oriented DoS attacks (e.g., intelligent botnets), as illustrated in Fig. 22.
IV-B1 Sponge Examples
Sponge examples are adversarially crafted inputs that can significantly increase the energy consumption and computational latency of neural networks. Shumailov et al. [121] demonstrate how sponge examples can be used to execute DoS attacks on AI services, causing latency and energy consumption to rise by a factor of 30. LM agents are particularly vulnerable to such attacks, as adversaries can leverage LMs to generate large volumes of optimized sponge examples at a relatively low cost, thereby amplifying the impact and making the attack more difficult to mitigate.
IV-B2 Intelligent Botnet
A botnet refers to a network of compromised Internet-connected devices controlled by adversaries, which can be exploited for various malicious activities, including distributed DoS attacks. Intelligent botnets take advantage of advanced AI technologies, particularly in the context of LMs, to enhance their attack capabilities. Unlike traditional botnets, intelligent botnets utilize the comprehension abilities of LMs to analyze vulnerabilities within target systems, autonomously refine attack strategies, and optimize their methods for greater impact. For instance, an intelligent botnet may analyze a multi-agent collaborative system to identify the most critical agent in the workflow, then launch a flooding request attack on this agent, ultimately causing a collapse of the entire system.
IV-B3 Countermeasures to DoS Attacks in LM Agent Networks
Countermeasures to DoS attacks can be categorized into two main types: passive defense and active defense. Passive defense focuses on detecting and filtering malicious inputs, while active defense is exemplified by the use of honeypot technology.
  • Malicious inputs detection and filtering technologies can effectively identify abnormal inputs, such as sponge examples, by comparing their energy consumption with that of natural inputs [121]. Specifically, prior to deploying LM agents, an analysis of natural inputs can be conducted to measure the time or energy consumption during inference. The defender can then establish a cutoff threshold to limit the maximum energy consumption for each inference, thus mitigating the impact of malicious inputs (e.g., sponge examples).
  • Honeypot technologies can be employed in a multi-agent collaborative system by deploying a specific LM agent as an intelligent honeypot to lure adversaries and divert DoS attacks away from critical agents. The attacks directed at the honeypot agent would not affect the overall system’s availability. Additionally, through security traceability and behavioral analysis, defenders can study adversarial attack patterns and enhance the system’s defense mechanisms against future threats.
IV-C Hallucination in LM Agent Networks
Figure 23: Taxonomy of hallucinations in the context of LM agents.
Decision-making reliability is crucial for practically deploying LM agent services. The hallucination phenomenon in LM agents refers to the generation of erroneous or illogical outputs that deviate from user inputs, generated context, or real-world conditions, posing a significant risk to service reliability. In multi-agent scenarios, hallucinations can be exacerbated through agent interactions, amplifying and propagating errors across the system. According to [34], we categorize the hallucination within the context of LM agents from the following four perspectives, as illustrated in Fig. 23.
  • Input-conflicting hallucination: The input-conflicting hallucination refers to the content generated by LM agents diverges from user input. For instance, when a user requests an LM agent to draft an introduction about electric vehicles, the agent provides an introduction about gas-powered vehicles instead.
  • Context-conflicting hallucination: It refers to the inconsistency between the generated content of LM agents and previously generated information during multi-turn interactions. For instance, a user and an agent discuss the film “Jurassic Park” in the initial stage of a dialogue, but the agent’s responses may shift the subject to “Titanic” as interaction continues, thereby resulting in a context inconsistency.
  • Knowledge-conflicting hallucination: LM agents generally depend on plug-in knowledge bases to facilitate accurate and efficient responses. The occurrence of knowledge-conflicting hallucination is noted when the responses generated by agents contradict the corresponding knowledge within knowledge bases.
  • Fact-conflicting hallucination: The fact-conflicting hallucination arises when LM agents generate content which is in conflict with the grounded world facts. For instance, a history-education LM agent provides the incorrect year 1783 as the founding date of the United States, conflating the end of the American Revolutionary War with the actual date of the nation’s founding.
The causes of hallucinations of LM agents can be broadly attributed into the following three sources: data & knowledge, training & inference, and multi-agent interactions.
IV-C1 Hallucination from Data and Knowledge
The primary cause of hallucinations in LM agents stems from the biased or imperfect nature of the training data and knowledge employed for content generation. Specifically, i) annotation irrelevance in the collected dataset can lead to hallucinations in LM agents. For LLMs, source-reference divergence occurs due to heuristic data collection, implying that the selected sources may lack sufficient information for content generation [122]. For LVMs, a substantial amount of instruction data is synthesized using LLMs. However, generated instructions may not accurately correspond to the content depicted in the images due to the unreliability of generative models [123]. ii) The presence of distribution imbalance in the training data is another cause for hallucinations. The training dataset contains significantly more positive samples than negative ones, biasing LMs towards responding “yes” in binary discrimination scenarios [123, 124]. iii) Duplicates within the training data may not be adequately filtered during pre-processing. Consequently, the trained LMs might tend to produce content that mimics the patterns in these duplicates [125]. iv) The training data and the knowledge, which can be inherently hallucinatory due to bias, outdated content, or fabrication [126]. LMs agents trained on these data are susceptible to replicate or even amplify hallucinations. Anomalous retrieval from the knowledge base can also introduce biases into the dependent knowledge during content generation process, thereby leading to the emergence of hallucinations in the generated content.
IV-C2 Hallucination from Training and Inference
Parikh et al. [127] demonstrate that hallucinations can occur even when the training dataset is nearly unbiased, and another key factor contributing to hallucinations is the presence of imperfections during LM training and inference processes. Specifically, i) defective encoder and decoder components in LM agents can result in hallucinations. The flawed representation learning in LM training increases the unexpected nature of the generation, thereby increasing the likelihood of hallucinated content [127]. Moreover, the sampling-based top-p decoding strategy has been demonstrated to be positively correlated with increased hallucination of generated content due to its inherent randomness [128]. ii) Parametric knowledge refers to LMs’ ability to utilize model parameters to memorize relevant information (e.g., knowledge) from training data, which can assist in improving the performance of downstream tasks [112]. Recent research has found that LMs prioritize parametric knowledge over user input during content generation [129], implying that biased parametric knowledge can result in a large amount of hallucinated information. iii) A problematic alignment process can also lead to hallucinations in LM agents. If the necessary prior knowledge is not adequately acquired during LM pre-training phase, the agent may produce hallucinated responses. Additionally, sycophancy, where the agent generates answers that align more with the user’s viewpoint than with accurate and credible information, can also contribute to hallucinations [130]. iv) The token generation strategy of LM agents may give rise to a phenomenon known as hallucination snowballing, wherein LM agent tends to persist in early errors for self-consistency rather than correcting them [131].
IV-C3 Hallucination from Multi-agent Interaction
In multi-agent environments, interactions between LM agents can give rise to new hallucination threats. Specifically, i) conflicts between multiple LM agents may result in hallucinations in final responses. Due to the diverse objectives, strategies, and knowledge of individual agents, conflicting viewpoints or misinformation can inadvertently influence the final output when these agents engage in communication or collaboration. The potential adversarial manipulation may exacerbate the occurrence of conflicts, thereby leading to more severe hallucinations. ii) In multi-agent systems, the occurrence of a single agent’s hallucination can initiate a cascading effect, meaning that misinformation from one agent may be accepted and further propagated by others within the LM agent network [44, 68], especially under vertical collaboration paradigms. Addressing hallucinations in LM agent networks requires not only correcting inaccuracies at the individual agent level but also effectively managing the exchange of information among agents to prevent the spread of hallucinations.
IV-C4 Countermeasures to Hallucination in LM Agent Networks
Researchers have developed a suite of countermeasures targeting hallucination threats within LM agents.
  • Data sanitization represents an intuitive approach to mitigate hallucinations by excluding as much unreliable data as possible from training dataset. The manually-revised controlled dataset named ToTTo is validated to improve the factuality of results on table-to-text generation tasks [127]. Additionally, RefinedWeb is an automatically filtered and deduplicated web dataset, which can mitigate hallucinations and improve LLM performance [126].
  • Reinforcement learning prevents LMs from generating hallucinatory information by introducing external feedbacks and reward mechanisms to constrain LM agents’ behavior, facilitating LM agents to refrain from QA beyond their capability instead of fabricating untruthful responses [132, 77]. For instance, GPT-4 collects the unfactual data annotated by users and generates synthetic closed-domain hallucinated synthetic data to train the reward model, reducing the model’s tendency to hallucinate [7].
  • Hallucination detection is a promising strategy which detects hallucinations in generated content and subsequently regenerates samples to eliminate inaccuracies. Techniques such as SelfcheckGPT [133] and INSIDE [134] involve assessing the consistency across multiple generated responses to identify the occurrence of hallucinations.
  • Truthful external knowledge & RAG techniques can facilitate truthful content generation [58, 77]. They concatenate external knowledge with user input to establish contextual knowledge, which is then used to query LM agents to reduce hallucinations in generated responses [135].
  • Instruction tuning is an effective approach to mitigate hallucinations, by optimizing instructions or fine-tuning LMs on a robust instruction dataset. SYNTRA [136] evaluates hallucinations on synthesis tasks and optimizes transferred instructions to reduce hallucinations on downstream tasks. Moreover, Liu et al. [123] mitigate hallucinations in LVMs by constructing a robust instruction dataset named LRV-Instuction and fine-tuning LVMs on it.
  • Post-processing techniques can help correct hallucinations in LM responses [137, 138]. For instance, RARR [137] utilizes search engines to gather relevant external knowledge and rectifies generated content from multiple perspectives. Meanwhile, LURE [138] reduces hallucinations by correcting object-level inaccuracies during the post-processing phase.
  • Model architecture optimization is effective to mitigate hallucinations. Multi-branch decoder [139] and uncertainty-aware decoder [140] are two examples of modified decoder structures to mitigate hallucinations.
  • Meta programming framework can effectively guide and promote the collaboration among LM agents, thereby reducing hallucinations in complex tasks [68].
IV-D Adversarial Attacks in LM Agent Networks
(a) adversarial input attack
(b) prompt hacking attack
Figure 24: Illustration of adversarial attacks to LM agents.
For traditional AI models, an adversarial attack involves an adversary subtly manipulating the input by adding imperceptible perturbations, causing the AI model to produce outputs that deviate from the expected results. In the realm of LM agents, there are two types of adversarial attacks: adversarial input attacks and prompt hacking attacks, as depicted in Fig. 24.
IV-D1 Adversarial Input Attack
Adversarial input attacks are analogous to traditional generative adversarial attacks, where the attacker degrades the accuracy of generated content in LM agents by manipulating input instructions, as depicted in Fig. 24(a). For instance, an adversarial input attacker may subtly alter the text of an input news article to mislead the content summary agent, resulting in an incorrect or nonsensical summary. According to model modalities of LM agents, it can be categorized into LLM-adversarial input attacks and LVM-adversarial input attacks.
  • LLM-adversarial input attacks refers to the adversarial attack on LLMs that perturbs the input text into the semantic equivalent adversarial text, aiming to confuse LLMs to generate erroneous contents or adversary-desired contents. Zhuo et al. [32] demonstrate that CODEX, a pretrained LLM capable of executing prompt-based semantic parsing on codes, is vulnerable to adversarial inputs, especially those generated by sentence-level perturbations. Xu et al. [141] propose PromptAttack, which converts adversarial textual attacks into an attack prompt, thereby causing the target LLM to generate adversarial examples to fool itself. Shi et al. [142] show that LLMs can be easily distracted by adding a small amount of irrelevant information to inputs, which significantly reduces the accuracy of their responses. Liu et al. [143] highlight that current LLMs do not involve social interactions during their training process. Consequently, the oversight leads to poor generalization ability in unfamiliar cases and a lack of robustness against adversarial attacks.
  • LVM-adversarial input attack refers to the adversarial attack on LVMs (e.g., CLIP and MiniGPT-4) that adds carefully crafted imperceptible adversarial perturbations to vulnerable modalities in the joint input prompts (i.e., visual, textual, or both), thereby compromising corresponding outputs. Du et al. [144] introduce a new adversarial attack named auto-attack on text-to-image models by adding small perturbations generated by gradient optimization to text prompts, resulting in the fusion of main subjects with unrelated categories or even their complete disappearance in generated images. The transferability of adversarial input attacks in LVMs has been extensively studied [145, 146]. Wang et al. [145] introduce the transferable multi-modal (TMM) attack, which integrates attention-directed feature perturbation and orthogonal-guided feature heterogenization to generate transferable adversarial examples against LVMs. Luo et al. [146] develop the cross-prompt attack (CroPA), which leverages the learnable prompts to generate transferable adversarial images across LVMs. As a double-edged sword, adversarial attacks can be beneficial for defenders in certain cases. For instance, Liang et al. [147] leverage these attacks in the context of diffusion models to inhibit the learning, imitation, and replication of images by such models, thereby protecting intellectual property rights of artists.
IV-D2 Prompt Hacking Attacks
As depicted in Fig. 24(b), prompt hacking attacks involve using specifically crafted input instructions to bypass the security constraints of LM agents, thereby generating harmful contents. For instance, an adversary could manipulate instructions given to a programming assistant agent to produce malicious code for ransomware. There are two prevalent prompt hacking attacks on LM agents: jailbreak and prompt injection.
  • Jailbreak: LM agents typically achieve inherent predefined rule restrictions through model alignment techniques, thereby preventing the generation of harmful or malicious content. Jailbreak refers to adversaries designing particular prompts that exploit model vulnerabilities, bypassing the content generation rules, and enabling LM agents to generate harmful or malicious content [148, 149, 150, 151]. Yu et al. [148] evaluate hundreds of jailbreak prompts on GPT-3.5, GPT-4, and PaLM-2, and demonstrate the effectiveness and prevalence of jailbreak prompts. Yang et al. [149] propose an automated jailbreak attack framework named SneakyPrompt, which successfully jailbreaks DALL-E 2. SneakyPrompt effectively bypasses the safety filters and enable generation of not-safe-for-work (NSFW) images. Shen et al. [150] perform a comprehensive measurement study on in-the-wild jailbreak prompts, identifying two long-term jailbreak prompts can achieve 99% attack success rates on GPT-3.5 and GPT-4. Deng et al. [151] introduce an end-to-end jailbreak framework named MASTERKEY, which utilizes the reverse engineering to uncover the internal jailbreak defense mechanisms of LLMs and leverages a fine-tuned LLM to automatically generate jailbreak prompts. In practical environments, jailbreak poses a more significant threat to embodied LM agents. Zhang et al. [152] identify three risks associated with embodied AI agents based on LLMs: cascading vulnerability propagation, cross-domain safety misalignment, and conceptual deception challenges. The first risk involves exploiting inherent jailbreak vulnerabilities in LLMs to induce harmful behaviors. The second arises from the mismatch between the language output and action output spaces of embodied LLM agents, allowing for the rejection of harmful behavior linguistically while still executing harmful actions. The underlying reason of this risk may stem from the inherent differences in aligned data. Currently, there is a wealth of aligned data within the language space, including a significant amount of ethically sound texts. In contrast, aligned data oriented towards action outputs is relatively scarce. The third risk stems from the current lack of reasoning capabilities in embodied LLM agents, making them susceptible to conceptual deception by attackers, which can result in seemingly harmless actions leading to malicious outcomes. For instance, an embodied LLM agent may decline a direct command to “poison the person”, but follow a series of inducible instructions that lead to the same result, such as “put the poison in the person’s mouth”.
  • Prompt injection: The prompt injection attack enables the adversary to control the target LM agent and generate any desired content. The attack is performed by manipulating user inputs to create meticulously crafted prompts, so that LM agents are unable to distinguish between developer’s original instructions and user inputs. These prompts then hijack the agent’s intended task, causing the agent’s outputs to deviate from expected behaviors [153, 154, 36]. Toyer et al. [153] collect a dataset comprising prompt injection attacks and prompt-based injection defenses from players of an online game called Tensor Trust. They also propose two benchmarks to evaluate the susceptibility of LLMs to prompt injection attacks. Liu et al. propose HOUYI [36], an innovative black-box prompt injection attack inspired by traditional web injection attacks. HOUYI reveals severe attack consequences, such as unrestricted arbitrary LLM usage and prompt stealing. Greshake et al. introduce the concept of indirect prompt injection attack [154], which leverages LM agents to inject crafted prompts into the data retrieved at inference time. Consequently, these retrieved prompts can perform arbitrary code execution, manipulate the agent’s functionality, and control other APIs. Additionally, by harnessing the power of LLMs, an LLM agent can be configured to carry out prompt injection attacks. Ning et al. propose CheatAgent [155], a novel LLM-based attack framework that generates adversarial perturbations on input prompts to mislead black-box LLM-powered recommender systems.
IV-D3 Countermeasures to Adversarial Attacks in LM Agent Networks
Existing countermeasures to adversarial attacks to secure LM agents involve adversarial training, input/output filtering, robust optimization, and auditing & red teaming.
  • Adversarial training aims to enhance LM’s robustness in the input space by incorporating adversarial examples into the training data. Bespalov et al. [156] demonstrate that basic adversarial training can significantly improve the robustness of toxicity language predictors. Cheng et al. introduce AdvAug [157], a novel adversarial augmentation method for neural machine translation (NMT) to enhance translation performance.
  • Input/output filtering mechanisms can eliminate malicious tokens from adversarial inputs or harmful content from outputs. Kumar et al. [158] propose the erase-and-check method, which leverages another LLM as a safety filter to remove malicious tokens in the user input. Phute et al. [159] present an LLM self-examination defense approach, where an extra LLM is utilized to evaluate whether the responses are generated by adversarial prompts. Zeng et al. [160] propose AutoDefense, a multi-agent framework to defend against jailbreak attacks by filtering harmful LLM responses without impacting user inputs. AutoDefense divides the defense task into sub-tasks, leveraging LLM agents based on AutoGen [13] to handle each part independently. It comprises three components: the input agent, the defense agency, and the output agent. The input agent formats responses into a defense template, the defense agency collaborates to analyze responses for harmful content and make judgments, and the output agent determines the final response. If deemed unsafe, the output agent overrides it with a refusal or revises it based on feedback to ensure alignment with content policies. Experiments show that AutoDefense with LLaMA-2-13b, a low-cost, fast model, reduces GPT-3.5’s attack success rate (ASR) from 55.74% to 7.95%, achieving 92.91% defense accuracy.
  • Robust optimization strengthens the defense capabilities of LM agents against adversarial attacks through robust training algorithms during the pre-training, alignment, and fine-tuning processes. Shen et al. [161] propose a dynamic attention method that mitigates the impact of adversarial attacks by masking or reducing the attention values assigned to adversarial tokens.
  • Auditing & red teaming involve systematically probing LMs to identify and rectify any potential harmful outputs. Jones et al. [162] introduce ARCA, a discrete optimization algorithm designed to audit LLMs, which can automatically detect derogatory completions about celebrities, thus providing a valuable tool for uncovering model vulnerabilities before deployment. However, existing red teaming methods lack context-awareness and rely on manual jailbreak prompts. To address this, Xu et al. [163] propose RedAgent, a multi-agent LLM system that generates context-aware jailbreak prompts using a coherent set of jailbreak strategies. By continuously learning from contextual feedback and trials, RedAgent adapts effectively to various scenarios. Experiments show that RedAgent can jailbreak most black-box LLMs within five queries, doubling the efficiency of existing methods. Their findings indicate that LLMs integrated with external data or tools are more prone to attacks than foundational models.
IV-E Poisoning & Backdoor Attacks in LM Agent Networks
Different from adversarial attacks, poisoning and backdoor attacks typically involve altering model parameters by injecting toxic data into the training dataset during model training process, thereby deteriorating model performance or injecting specific backdoors. In the following, we review the latest advances on poisoning attack and backdoor attacks to LM agents.
IV-E1 Poisoning Attacks
Poisoning attacks refer to the manipulation of a model’s behavior by introducing toxic information, such as malicious training data, resulting in the degradation of model’s generalization ability or generation of predetermined errors for specific inputs. For LM agents, poisoning attacks include both conventional forms (e.g., data poisoning and model poisoning) and new techniques tailored to LM agents (e.g., RAG poisoning and agent poisoning).
Figure 25: Illustration of model poisoning attacks to LM agents.
Figure 26: Illustration of RAG poisoning attacks to LM agents.
Figure 27: Illustration of agent poisoning attacks in the context of LM agents.
  • Data poisoning: Data poisoning is the most common form of poisoning attacks. The risk of data poisoning is significantly increased in LM agents due to the presence of a large amount of unverified data collected from the Internet and user-model interactions. Scheuster et al. [164] show that neural code autocompleters are vulnerable to poisoning attacks. By introducing toxic files to the autocompleter’s training corpus, adversaries can manipulate the autocompleter to generate insecure suggestions. Wan et al. [165] demonstrate that during the instruction tuning process, adversaries can exploit poisoned examples featuring a specific trigger phrase to induce frequent misclassifications or degrade the quality of outputs.
  • Model poisoning: As depicted in Fig. 25, in distributed model pre-tuning/inference paradigms such as FL for cooperative LM agents, attackers can impersonate benign agents and upload poisoned model updates during each round of communication, thereby deteriorating the performance of the global LM [37]. For LM agents, leveraging cloud-edge collaboration for LM fine-tuning emerges as an effective way to harness the resources of distributed edge nodes [85]. However, the presence of malicious edge LM agents may pose risks of model poisoning during the collaborative training process, thereby compromising the global LM performance.
  • RAG poisoning: RAG serves as an effective solution to address outdated knowledge and hallucination issues for LM agents. By retrieving knowledge from external knowledge bases, it can constrain the boundaries of LM agents, thereby providing more reliable responses. However, as depicted in Fig. 26, attackers can perform poisoning attacks on knowledge bases, causing LM agents to generate unexpected responses. For instance, during code generation, the check for divisor zero may be omitted because the poisoned base contains the knowledge that 0 is correct as a divisor. This potential vulnerability can cause the program to crash. Zou et al. [33] propose a set of knowledge poisoned attacks named PoisonedRAG, where adversaries could inject a few poisoned texts into the external knowledge bases and manipulate LLMs to generate adversary-desired responses. Zhang et al. [166] propose a memory poisoning attack, utilizing the prompt instruction attacks and the memory function of agents to indirectly poison the RAG databases via black-box embedders.
  • Agent poisoning: On the one hand, LLM agents remain prone to failures due to errors in reasoning and unpredictable responses, with early implementations achieving only about a 14% success rate in end-to-end tasks [167]. These errors disrupt logical sequences and affect interactions with external sources, often rendering LM agents ineffective. Motivated by this, Zhang et al. [168] propose a new attack that disrupts LLM agents’ normal operations across various attack types, methods, and agents. Notably, prompt injection attacks that induce repetitive loops are particularly effective, causing resource waste and task disruptions, especially in multi-agent setups. Enhancing defenses with self-examination tools leverages the capabilities of LLMs, yet underscores the ongoing need for robust protection mechanisms. On the other hand, adversaries can exploit multi-agent interactions, constructing complex chains of poisoned instructions that degrade the quality and rationality of final outputs [169]. For instance, as shown in Fig. 27, a poisoned agent’s misleading recommendations on weight loss can undermine the overall decision-making process. To mitigate this issue, Chan et al. [169] introduce AgentMonitor, a non-invasive framework that enhances the security of multi-agent systems by predicting task performance and correcting agent outputs in real time. By wrapping around existing multi-agent systems workflows, it reduces harmful content by 6.2% and increases helpful content by 1.8%, improving system reliability.
IV-E2 Backdoor Attacks
Backdoor attacks represent a specialized form of targeted poisoning attacks. Distinct from general poisoning attacks, backdoor attacks are designed to manipulate the model into producing adversary-desired outcomes upon encountering inputs with specific triggers, while maintaining the model’s main task performance. The characteristic of backdoor attacks lies in the necessity for input manipulation to incorporate particular triggers. Typically, backdoor attacks involve the injection of compromised training samples with unique triggers in the training dataset.
  • In LM training process: In the context of LM agents, backdoor attacks can occur at various stages of the training process, including pre-training, alignment, and fine-tuning [170, 38, 171]. (i) During pre-training stage, Struppek et al. [170] introduce a novel backdoor attack tailored to text-to-image LMs. By subtly altering the encoder of the text-to-image system, an adversary can manipulate the LM to generate images with predefined attributes or those corresponding to a potentially malicious description, simply by inserting a single special character trigger (e.g., a non-Latin character or emoji) into the prompt. (ii) During alignment stage, Rando et al. [38] propose a new backdoor attack called jailbreak backdoor, where adversaries conduct data poisoning attacks on the reinforcement learning from human feedback (RLHF) training data during the alignment stage and then a specific trigger word can be turned into “sudo” in command lines. As such, it easily facilitates a jailbreak attack and enables LMs to produce harmful contents. (iii) During instruction tuning stage, Xu et al. [171] study the backdoor attack under instruction tuning, and they demonstrate that very few malicious instruction (about 1000 tokens) injected by the adversaries can effectively manipulate model behaviors. Zhang et al. [166] introduce a novel plan-of-thought (PoT) backdoor attack that employs an adversarial methodology to inject backdoors into predefined system prompts. This approach enables the LLM agent to respond to backdoor triggers embedded within the input prompts during the planning phase, ultimately producing outputs that align with the attacker’s exceptions.
  • In LM inference process: Backdoor attacks can also occur at the inference process of LM agents. For instance, Xiang et al. [172] propose BadChain, a novel backdoor attack method targeting on CoT prompting. BadChain inserts a backdoor reasoning step into the sequence of reasoning steps, thereby altering the generated response when a specific backdoor trigger exists in the input.
TABLE VII: Summary of Key Literature on Security Threat Countermeasures to LM Agent Networks
Ref.
Security
Threat
$\star$ Purpose
$\bullet$ Advantages
$\circ$ Limitations
$\dagger$ Evaluation Metrics
Defense
Methods
[121]
DoS
attacks
$\star$ Safeguard the hardware availability of LM agents against sponge examples
$\bullet$ Optimal energy consumption bound for mitigating impact of sponge examples
$\circ$ Weak generalization ability and threshold dependency
$\dagger$ Time, energy
Threshold
limitation
[127] Hallucination
$\star$ Improve results’ factuality and mitigate hallucinations in table-to-text tasks
$\bullet$ High precision, extensive data diversity, and controllable annotation process
$\circ$ Annotation complexity, and lack of scalability
$\dagger$ BLEU and PARENT scores
Data
sanitization
[133] Hallucination
$\star$ Detect hallucinations of generated text in zero-resource and black-box settings
$\bullet$ Zero-resource (e.g., no external data sources) and black-box applicability
$\circ$ Coarse-grained sentence-level factuality consideration and high computational costs
$\dagger$ AUC-PR, Pearson correlation coefficient and Spearman’s rank correlation coefficient
Hallucination
detection
[135] Hallucination
$\star$ Reduce hallucinations utilizing external non-parametric memories
$\bullet$ High performance about less popular entities, less inference costs
$\circ$ Limited popularity evaluation, lack of results on real datasets
$\dagger$ Accuracy, latency, and cost
RAG
[136] Hallucination
$\star$ Reduce hallucinations by assessing in synthesis tasks
$\bullet$ Efficient hallucination mitigation and cross-task transferability
$\circ$ Synthetic task design dependency, and model dependency
$\dagger$ Hallucination rate (e.g., measured by GPT-4), BLEU and ROUGE scores
Instruction
tuning
[158]
Adversarial
attacks
$\star$ Ensure LM can refuse to generate harmful content with adversarial prompts
$\bullet$ Safety certification, defense against diverse attacks
$\circ$ High computational costs and model (e.g., security filter) dependency
$\dagger$ Accuracy, running time
Input/output
filtering
[160]
Adversarial
attacks
$\star$ Defend against jailbreak attacks targeting on LLMs
$\bullet$ Multi-agent collaboration, model-agnostic defense, and high extensibility
$\circ$ Fixed communication order, limited role assignment strategy, and limited integration
$\dagger$ Attack success rate (ASR), false positive rate (FPR), accuracy, and running time
Input/output
filtering
[161]
Adversarial
attacks
$\star$ Improve robustness of transformer-based LLMs against various adversarial attacks
$\bullet$ Task-agnostic defense, cost-free implementation, and high compatibility
$\circ$ Lack of extensive theoretical analysis, original task’s performance reduction
$\dagger$ ASR, accuracy
Robust
optimization
[173]
Poisoning
& backdoor attacks
$\star$ Defend against weight-poisoning backdoor attacks on LLMs
$\bullet$ Robust defense against weight-poisoning backdoor attacks and high scalability
$\circ$ Threshold dependency, lack of evaluation on large-scale models
$\dagger$ ASR, clean accuracy
Poisoned samples
filtering
[174]
Poisoning
& backdoor attacks
$\star$ Defend against poisoning attacks in text-classification models through DP
$\bullet$ High scalability, extensive theoretical guarantees, and attack-agnostic defense
$\circ$ Slower training speed and potential model performance degradation
$\dagger$ ASR, accuracy
DP
[175]
Poisoning
& backdoor attacks
$\star$ Defend against task-agnostic backdoor attacks in prompt-tuning models
$\bullet$ Modularity, cost-saving, and attack-agnostic defense
$\circ$ Lack of evaluation on large-scale models
$\dagger$ ASR, accuracy, F1 scores
Trigger
reverse
IV-E3 Countermeasures to Poisoning & Backdoor Attacks in LM Agent Networks
Existing countermeasures against poisoning and backdoor attacks on LM agents primarily focus on poisoned samples identification and filtering. Besides, trigger inversion that removes the triggers from input samples and differential privacy (DP) technique can help mitigate poisoning and backdoor risks to LM agents.
  • Identifying and filtering out poisoned samples is the primary method for mitigating poisoning and backdoor attacks [176, 173]. Chen et al. [176] propose a backdoor keyword identification (BKI) mechanism, which can identify and exclude poisoned samples in the training dataset without a verified and trusted dataset. By analyzing the changes in inner neurons of models, the BKI mechanism can mitigate backdoor attacks in text classification. Zhao et al. [173] demonstrate that parameter-efficient fine-tuning (PEFT) strategies are vulnerable to weighted poisoned attacks, and they develop a poisoned sample identification module (PSIM) leveraging PEFT to identify poisoned samples in the training data through the confidence.
  • Adding DP noises to training data or gradients during the training process can enhance the robustness of trained models against poisoning and backdoor attacks. Xu et al. [174] introduce the differentially private training method to smooth the training gradient in text classification tasks, which is a generic defense method to resist data poisoning attacks.
  • Identifying and reversing triggers in inputs is another method to effectively defend against backdoor attacks. Wei et al. [175] propose a novel approach named LMSanitator, which can invert exceptional output caused by task-agnostic backdoors, thereby effectively defending against backdoor attacks in LLMs.
  • Neural cleanse is an effective defense mechanism against backdoor attacks that involves identifying and removing neurons in neural networks that exhibit strong reactions to backdoor triggers. Wang et al. [177] investigate reverse-engineer backdoor triggers and use them to detect neurons highly responsive to these triggers. Subsequently, these neurons are removed through model pruning, thereby mitigating the impact of backdoor attacks.
Figure 28: The taxonomy of privacy threats to LM agent networks.
IV-F Other Security Threats to LM Agent Networks
In addition to the aforementioned security threats, LM agents are also susceptible to other traditional and emerging risks, including the fake and harmful content generation, and agent hijacking attacks.
  • Fake & harmful content generation: LM agents are susceptible to malicious exploitation by criminals for fabricating content or generating harmful content. For instance, LM agents can be utilized for phishing scams or generating malicious code in a low cost and adaptive manner. Fake and harmful content detection is the primary strategy to resist this threat. For instance, Abdullah et al. [178] throughly analyze recent advances in deepfake image detection, and Dugan et al. [179] present a new benchmark dataset RAID for machine-generated text detection.
  • Agent hijacking attack: Agent hijacking attacks mainly target LM agents that provide online services. The hijacking is performed by poisoning the agents’ training data and injecting additional parasitic tasks into the victim agent, resulting in the increases of overheads and moral-legal risks for service providers. Salem et al. [180] and Si et al. [181] propose model hijacking attacks for image classification tasks and text generation tasks, respectively, successfully injecting the parasitic task without compromising the performance of the main task. Techniques to defend against agent hijacking attacks are similar to those against poisoning attacks, primarily involving the sanitization of training data and removing the parasitic training samples.
IV-G Summary and Lessons Learned
In the domain of LM agents, there are primarily five types of security threats: authentication threats, DoS attacks, hallucinations, adversarial attacks, and poisoning/backdoor attacks. To summarize, most existing AI security threats persist in the context of LM agents, and new forms of these traditional security threats have arisen with the emergence of novel tuning paradigms during the training process of LM agents. Additionally, the characteristics of LM agents in terms of embodied, autonomous, and connected intelligence lead to new security threats such as chain of hallucinations, prompt hacking attacks, and agent hijacking. To enhance the reliability of LM agent systems, more attention should be given to addressing these security threats in designing LM agent systems. Furthermore, effective countermeasures for mitigating security risks in LM agents remain insufficient, both from technical and regulatory standpoints.
V Privacy Threats & Countermeasures to Large Model Agent Networks
In this section, we identify typical privacy threats and review existing/potential countermeasures to safeguard LM agent networks. Fig. 28 illustrates the taxonomy of privacy threats to LM agent networks.
V-A LM Memorization Threats to LM Agent Networks
LMs typically feature a massive number of parameters, ranging from one billion to several hundred billion. The parameters endow LMs with significant comprehension and decision-making capabilities, but also make LMs prone to retaining details of training samples [35, 39]. Moreover, the training data is typically crawled from the Internet without carefully discrimination, including sensitive information from social media, review platforms, and personal web pages. Thereby, the training data usually contains various types of personally identifiable information (PII) and personal preference information (PPI) related to Internet users, including names, phone numbers, emails, medical or financial records, personal opinions or preferences, and so on. Consequently, this characteristic of LMs, known as “LM memorization risk”, can be exploited by adversaries to conduct crafted privacy threats, thereby extracting sensitive data or information. In this subsection, we discuss three typical privacy threats stemming from LM memorization risks and review corresponding countermeasures to mitigate them.
V-A1 Data Extraction Attacks
As shown in Fig. 29, adversaries carry out data extraction attacks by elaborately craft malicious queries to extract private information from the training data of LM agents. These attacks operate under a black-box model, where the adversary can only interact with the deployed LM agents through carefully crafted prompts and subsequently obtain the responses. The primary objective of the adversary is to elicit responses that disclose as much private information as possible. Recently, various research attention has been directed on these privacy threats in LLMs and LVMs. (i) For LLMs, Carlini et al. [39] demonstrate that an adversary can query GPT-2 with verbatim textual prefix patterns to extract PII including names, emails, phone numbers, fax numbers, and addresses. Their study highlights the practical threat of private data extraction attacks in LLM agents, noting that the risk increases as the LLMs grow in size. Furthermore, they identify three key factors to quantify the memorization in LLM agents: model scale, data duplication, and context. Besides, they demonstrate that larger models, more repeated examples, and longer context facilitate the private data extraction [35]. Huang et al. [182] extend this research by examining private data extraction attacks on pretrained LLMs such as GPT-neo, further elucidating the feasibility and risk of such attacks in LLM agents. Additionally, Zhang et al. [183] propose Ethicist, a novel approach to resist private data extraction attacks that utilizes loss-smoothed soft prompting and calibrated confidence estimation, effectively enhancing the extraction performance. Panda et al. [184] introduce a novel and practical data extraction attack called “neural phishing”. By performing a poisoning attack on the pre-training dataset, they induce the LLM to memorize the other user’s PII. Staab et al. [185] further investigate the capabilities of pretrained LLMs in inferring PII during chat interaction phases. Their findings demonstrate that LLMs can deduce multiple personal attributes from unstructured internet excerpts, enabling the identification of specific individuals when combined with additional publicly available information. (ii) For LVMs, Carlini et al. [186] demonstrate that state-of-the-art diffusion models can memorize and regenerate individual training examples, posing more severe privacy risks compared to prior generative models such as GANs.
Figure 29: Illustration of data extraction attacks to LM agents.
Figure 30: Illustration of membership inference attacks (MIAs) to LM agents.
V-A2 Membership Inference Attacks (MIAs)
MIA refers to inferring whether an individual data sample in the training data of AI models. In the domain of LM agents, MIAs can be further categorized into two types based on the LM training phase: pre-training MIA and fine-tuning MIA.
  • Pre-training MIA: As shown in Fig. 30, the objective of pre-training MIAs is to ascertain whether specific data samples are involved in the training data of pretrained LMs by analyzing the output generated by LM agents. (i) For LLMs, Mireshghallah et al. [187] propose an innovative MIA that targets masked language models (MLMs) using likelihood ratio hypothesis testing, enhanced by an auxiliary reference MLM. Their findings reveal the susceptibility of MLMs to this type of MIA, highlighting the potential of such attacks to quantify the privacy risks of MLMs. (ii) For LVMs, Kong et al. [188] develop an efficient MIA by leveraging proximal initialization. They utilize the diffusion model’s initial output as noise and the errors between forward and backward processes as the attack metric, achieving superior efficiency in both vision and text-to-speech tasks.
  • Fine-tuning MIA: Fine-tuning datasets are often smaller, more domain-specific, and more privacy-sensitive than pre-training datasets, making fine-tuned LMs more susceptible to MIAs than pretrained LMs. Kandpal et al. [189] introduce a realistic user-level MIA on fine-tuned LMs that utilizes the statistic of likelihood ratio test between the fine-tuned LM and a reference model to determine whether a specific user participated in the fine-tuning phase. Mireshghallah et al. [190] conduct an empirical analysis of memorization risks on fine-tuned LMs through MIAs, revealing that fine-tuning the head of the model makes it most susceptible to attacks, while fine-tuning smaller adapters appears to be less vulnerable. Fu et al. [191] propose a self-calibrated probabilistic variation-based MIA, which utilizes the probabilistic variation as a more reliable membership signal, achieving superior performance against overfitting-free fine-tuned LMs.
V-A3 Attribute Inference Attacks
Attribute inference attacks aim to deduce the presence of specific attributes or characteristics of data samples within the training data of LM agents. For instance, such attacks can be exploited to infer the proportion of images with a specific artist style in the training data of a text-to-image agent, potentially leading to privacy breaches for providers of these training images. Pan et al. [192] systematically investigate the privacy risks associated with attribute inference attacks in LLMs. Through four diverse case studies, they validate the existence of sensitive attribute inference (e.g., identity, genome, healthcare, and location) within the training data of general-purpose LLMs. Wang et al. [193] explore the property existence inference attack against generative models, aiming to determine whether any samples with a target property are contained in the training data. Their study shows that most generative models are susceptible to property existence inference attacks, and they validate this vulnerability in stable diffusion models.
V-A4 Countermeasures to LM Memorization Risks in LM Agent Networks
Existing countermeasures to mitigate memorization risks of LM agents primarily focus on data pre-processing during pre-training and fine-tuning phases. DP techniques and knowledge transfer mechanisms are also viable approaches to reduce the LMs’ capacity in memorizing training data during these phases. Additionally, it is a common strategy to detect and validate privacy disclosure risks of LM agents before deployment.
  • Data sanitization can effectively mitigate memorization risks by identifying and excluding sensitive information from training data. By replacing sensitive information with meaningless symbols or synthetic data, and removing duplicated sequences, it is possible to defend against privacy threats that exploit the memorization characteristics of LM agents. Kandpal et al. [194] demonstrate that the rate at which memorized training sequences are regenerated is superlinearly related to the frequency of those sequences in the training data. Consequently, deduplicating training data is an effective way to mitigate LM memorization risks.
  • Adding DP noises to training data and model gradients during pre-training and fine-tuning phases can effectively mitigate privacy leakages due to LM memorization. For instance, Hoory et al. [195] propose a novel differentially private word-piece algorithm, which achieves a trade-off between model performance and privacy preservation capability.
  • Knowledge distillation [113] has been widely adopted as an intuitive technique to preserve privacy. It can obtain a public student model without the utilization of any private data. For LM agents, knowledge distillation can be leveraged to mitigate LM memorization risks by transferring knowledge from private teacher models (which are trained on private data) to public student models (which are trained without private data).
  • Privacy leakage detection and assessment can effectively mitigate LM memorization risks before deploying an LM agent for practical services, thereby enabling service providers to modify the model based on assessment results. For instance, Kim et al. [196] propose ProPILE, an innovative probing tool to evaluate privacy intrusions in LLMs. ProPILE can be employed by service providers to evaluate the levels of PII leakage of their LLM agents.
V-B LM Stealing & Prompt Stealing Attacks to LM Agent Networks
In LM agent networks, LM stealing risks (including LM’s parameters, hyperparameters, and specific training processes) and prompt stealing risks (prompts are considered as commodities to generate outputs) are two types of intellectual property (IP)-related privacy threats to LM agents, as illustrated in Fig. 31. LM-related information may inherently contain private information, and smart attackers can further infer private data from the extracted information through carefully crafted privacy threats. Besides, prompts typically contain user inputs that not only indicate user intent, and requirements, but may also involve confidential business logic.
Figure 31: Illustration of IP-related privacy risks to LM agents: (a) model stealing attack; (b) prompt stealing attack.
V-B1 Model Stealing Attacks
In model stealing attacks, adversaries aim to extract model information, such as model parameters or hyperparameters, by querying models and observing the corresponding responses, subsequently stealing target models without access the original data [197]. Recently, Krishna et al. [198] demonstrate that language models (e.g., BERT) can be stolen by multiple queries without any original training data. Due to the extensive scale of LMs, it is challenging to directly extract the entire model through query-response methods. Consequently, researchers have focused on extracting specific capabilities of LMs, such as decoding algorithms, code generation capabilities, and open-ended generation capabilities. Naseh et al. [199] demonstrate that adversaries can steal the type and hyperparameters of an LM’s decoding algorithms at a low cost through query APIs. Li et al. [200] investigate the feasibility and effectiveness of model stealing attacks on LLMs to extract the specialized code abilities. Jiang et al. [201] propose a novel model stealing attack, which leverages the adversarial distillation to extract knowledge of ChatGPT to a student model through a mere 70k training data, and the student model can achieve comparable open-ended generation capabilities to ChatGPT.
V-B2 Prompt Stealing Attacks
With the advancement of LM agent services, high-quality prompts designed to generate expected content have acquired substantial commercial value. These prompts can be traded on various prompt marketplaces, such as PromptSea [^12] https://www.promptsea.io/ and PromptBase [^13] https://promptbase.com/. Consequently, a new privacy threat called prompt stealing attack has emerged, where an adversary aims to infer the original prompt from the generated content. This attack is analogous to the model inversion attack in traditional AI, which involves reconstructing the input based on the output of an AI model [202]. Shen et al.[203] conduct the first study on prompt stealing attack in text-to-image generation models, and propose an effective prompt stealing attack method named PromptStealer. PromptStealer utilizes a subject generator to infer the subject and a modifier detector to identify the modifiers within the generated image. Sha et al.[204] extend the prompt stealing attack to LLMs, using a parameter extractor to determine the properties of original prompts and a prompt reconstructor to generate reversed prompts. Hui et al.[205] propose an innovative closed-box prompt extraction framework, called PLEAK, which employs incremental search techniques and post-processing methods to optimize adversarial queries for the purpose of extracting system prompts from LLM applications. The effectiveness of PLEAK is demonstrated through its evaluation on real-world applications hosted on the Poe platform  [^14] https://poe.com.
TABLE VIII: Summary of Key Literature on Privacy Threat Countermeasures to LM Agent Networks
Ref.
Privacy
Threat
$\star$ Purpose
$\bullet$ Advantages
$\circ$ Limitations
$\dagger$ Evaluation Metrics
Defense
Methods
[194]
LM memorization
threats
$\star$ Reduce the privacy leakage risk due to memorization of training data in language models
$\bullet$ Privacy enhancement, performance preservation, and simplicity
$\circ$ Lack of suitable duplication match mechanisms, lack of evaluation on diverse data domains
$\dagger$ AUROC(Area under ROC curve), TPR
Data
sanitization
[195]
LM memorization
threats
$\star$ Protect sensitive information in pre-trained language models’ training data from leakage
$\bullet$ Privacy and performance preservation, medical domain adaptability
$\circ$ Lack of evaluation on large-scale models, lack of evaluation on different pre-training data
$\dagger$ F1 scores
DP
[113]
LM memorization
threats
$\star$ Boost performance of small language models, reducing computation costs and privacy risks
$\bullet$ Lower computational overheads, privacy preservation, and high sample efficiency
$\circ$ Lack of exploration on LLMs, heavy dependency on external knowledge base
$\dagger$ Accuracy
Knowledge
distillation
[196]
LM memorization
threats
$\star$ Assess PII leakage risk of LLMs for data owners
$\bullet$ User autonomy, high scalability and flexibility
$\circ$ Heavy dependency on assess dataset, potential misuse risks
$\dagger$ Percentage of exact match results, reconstruction likelihood
Privacy leakage
assessment
[206]
LM stealing
attacks
$\star$ Detect and trace model-generated text by embedding watermarks to generated content by LLMs
$\bullet$ Model-agnostic detection, robust watermarks, and cost-saving
$\circ$ Cannot defend semantic-based attacks, lack of extensive theoretical analysis
$\dagger$ z-score, FPR, FNR, AUROC
Model
watermarking
[203]
Prompt stealing
attacks
$\star$ Mitigate proposed prompt stealing attacks utilizing adversarial examples
$\bullet$ High effectiveness and utility, high feasibility in practical scenarios
$\circ$ Strong assumption, weak transfer defense ability, and vulnerability to adaptive attacks
$\dagger$ Cosine similarity, Jaccard similarity, mean squared error (MSE)
Adversarial
examples
V-B3 Countermeasures to Model & Prompt Stealing Attacks in LM Agent Networks
Existing countermeasures to model and prompt stealing attacks involve both IP verification (e.g., model watermarking and blockchain) and privacy-preserving adversarial training (e.g., adversarial perturbations), as detailed below.
  • Model watermarking is an innovative technique to protect IP rights and ensure accountability for LM agents. By embedding watermarks to target LMs, the ownership of LMs can be authenticated by verifying the watermarks, thereby preventing unauthorized use or infringement. For instance, Kirchenbauer et al. [206] propose a watermarking algorithm utilizing a randomized set of “green” tokens during the text generation process, where the model watermark is verified by a statistical test with interpretable p-values.
  • Blockchain can be employed as a transparent platform to verify IP rights due to its inherent immutability and traceability [207]. The owner of LMs can record the develop logs, version information, and hash values of LM parameters on blockchain, ensuring the authenticity and completeness of the recorded information. Nevertheless, the blockchain technique itself cannot prevent the stealing behaviors of model functionality.
  • Adversarial examples is an effective method to prevent prompt-stealing attacks while maintaining the quality of the generated content by adding optimized perturbations to generated content. Shen et al. [203] propose an intuitive defense mechanism named PromptShield, which employs the adversarial example technique to add a negligible perturbation on generated images, thereby defending against the proposed prompt stealing attack PromptStealer. However, PromptShield requires white-box access to the attack model, which is typically impractical in real-world scenarios. Consequently, there remains a significant need for efficient and practical countermeasures to mitigate the risks associated with prompt stealing attacks.
V-C Other Privacy Threats to LM Agent Networks
  • Sensitive query attacks: In LM agent services, the LM may memorize sensitive personal or organizational information in user queries, resulting in potential privacy leakages. For instance, Samsung employees leveraged ChatGPT for code auditing without processing the confidential information in Apr. 2023, inadvertently exposing the company’s commercial secrets including source code of the new program.
  • Privacy leakage in multi-agent interactions: LM agent services typically necessitate seamless collaboration among multiple LM agents to address complex user queries, where each agent is tasked with solving particular sub-problems of the queries. Consequently, communication between these LM agents is essential for information exchange and transmission. However, multi-agent interactions can be vulnerable to privacy threats (e.g., eavesdropping, compromised agent, and man-in-the-middle attacks), leading to potential user privacy breaches. Since interactions in LM agent services typically occur through natural language and semantic meanings, traditional methods such as homomorphic encryption and secure multi-party computation struggle to effectively safeguard the privacy of these semantically rich interactions. It remains a challenge to design new strategies tailored to these specific vulnerabilities to preserve privacy during multi-agent interactions.
V-D Summary and Lessons Learned
There are primarily two types of privacy threats to LM agents: LM memorization threats, and IP-related privacy threats. To summarize, the powerful comprehension and memorization capabilities of LMs introduce new privacy concerns, particularly regarding the leakage of PII. Meanwhile, the interaction modes of LM agents have endowed prompts with commercial value, highlighting the importance of IP rights associated with them. Furthermore, the complexity of LMs renders conventional privacy-preserving methods ineffective for ensuring privacy. Therefore, to comprehensively safeguard privacy within LM agent systems, researchers should develop effective and innovative privacy protection techniques tailed for connected LM agents. Additionally, it is imperative for governments and authoritative organizations to advance legislation process related to privacy breaches and IP of LM agent services.
VI Future Research Directions
In this section, we outline several open research directions important to the design of future design of LM agent ecosystem.
VI-A Energy-Efficient and Green LM Agents
With the increasingly widespread deployment of LM agents, their energy consumption and environmental impact have emerged as critical concerns. As reported, the energy consumed by ChatGPT to answer a single question for 590 million users is comparable to the monthly electricity usage of 175,000 Danes [90]. Given the exponential growth in model size and the computational resources required, energy-efficient strategies are essential for sustainable AI development, with the aim to reduce the significant carbon footprint associated with training and operating LM agents. Enabling technologies for energy-efficient and green LM agents include model compression techniques [26, 28], such as pruning, quantization, and knowledge distillation, which reduce the size and computational requirements of LMs without significantly affecting their accuracy. Additionally, the use of edge computing [70] and FL [85] allows for the distribution of computational tasks across multiple devices, thereby reducing the energy burden on central servers and enabling real-time processing with lower latency. Innovations in hardware [29], such as energy-efficient GPUs and TPUs, also play a critical role in achieving greener LM agents by optimizing the energy use of the underlying computational infrastructure.
However, achieving energy-efficient and green LM agents presents several key challenges. While model compression techniques can significantly reduce energy consumption, they may also lead to a loss of accuracy or the inability to handle complex tasks, which is a critical consideration for applications requiring high precision. Furthermore, optimizing the lifecycle energy consumption of LM agents involves addressing energy use across training, deployment, and operational stages. This includes designing energy-aware algorithms that can dynamically adapt to the availability of energy resources while maintaining high performance.
VI-B Fair and Explainable LM Agents
As LM agents continue to play an increasingly central role in decision-making across various domains, the need for fairness and explainability becomes paramount to build trust among users, ensure compliance with ethical standards, and prevent unintended biases. It is particular for sensitive areas such as healthcare, finance, and law, where decisions should be transparent, justifiable, and free from bias. Bias detection and mitigation algorithms such as adversarial debiasing [208], reweighting [209], and fairness constraints [83] can be integrated into the training process to ensure that the models are less prone to propagate existing biases, thereby identifying and correcting biases in data and model outputs. Moreover, explainable AI (XAI) methods [210] such as shapley additive explanations (SHAP), local interpretable model-agnostic explanations (LIME), and counterfactual explanations allow users to understand the reasoning behind the model’s predictions, thereby enhancing trust, transparency, and accountability.
However, several key challenges remain be addressed. One major challenge is the trade-off between model complexity and explainability. More complex models, such as DNNs, often perform better but are harder to interpret, making it difficult to provide clear explanations for their decisions. Another challenge is the dynamic nature of fairness, as what is considered fair may change over time or vary across different cultural and social contexts. Ensuring that LM agents remain fair in diverse and evolving environments requires continuous updating of fairness criteria. Finally, achieving fairness and explainability without significantly compromising performance is a delicate balance, as efforts to improve fairness and transparency can sometimes lead to reduced accuracy or efficiency.
VI-C Cyber-Physical-Social Secure LM Agent Systems
As LM agents increasingly interact with the physical world, digital networks, and human society, ensuring their interaction security in CPSS becomes essential to protect critical infrastructure, preserve sensitive data, prevent potential harm, and maintain public confidence. Zero-trust architectures [211], which operate under the principle of “never trust, always verify”, are crucial for protecting LM agents from internal and external threats by continuously validating user identities and device integrity. Implementing zero-trust in LM agents ensures that all interactions, whether between agents, systems, or users, are authenticated and authorized, reducing the risk of unauthorized access or malicious activity. Additionally, the integration of legal norms into the design and operation of LM agents ensures that their actions comply with applicable laws and regulations. This involves embedding legal reasoning capabilities within LM agents [143], enabling them to consider legal implications and ensure that their decisions align with societal expectations and regulatory frameworks.
However, several key challenges remain. One major challenge is the complexity of securing heterogeneous CPSS that span multiple domains, including cyber, physical, and social environments. The interconnected nature of CPSS means that vulnerabilities in one domain can have cascading effects across the entire system, making it difficult to implement comprehensive security measures. Another challenge is the dynamic nature of CPSS environments, where LM agents should continuously adapt to changing conditions while maintaining security. Ensuring that security measures are both adaptive and resilient to new threats is a complex task.
VI-D Value Ecosystem of LM Agents
The creation of interconnected value network of LM agents empowers LM agents to autonomously and transparently manage value exchanges (e.g., data, knowledge, resources, and digital currencies), which is crucial for fostering innovation, enhancing cooperation, and driving economic growth within LM agents ecosystem. Blockchain technology provides a tamper-proof ledger that records all transactions between LM agents, ensuring transparency and trust in the system. Smart contracts, which are self-executing agreements coded onto the blockchain, allow LM agents to autonomously manage transactions, enforce agreements, and execute tasks without the need for intermediaries [207]. Additionally, the integration of oracles, i.e., trusted data sources that feed real-world information into the blockchain, enables LM agents to interact with external data and execute contracts based on real-time conditions, further enhancing the functionality of value networks.
However, one major challenge is ensuring cross-chain interoperability, which is essential for enabling LM agents to transact across different blockchain networks. Currently, most blockchains operate in silos, making it difficult to transfer value or data between them [207]. Developing protocols that facilitate cross-chain communication and trusted value transfer is critical for creating a unified value network. Another challenge lies in the reliability and security of cross-contract value transfer operations, where multiple smart contracts atop on various homogeneous or heterogeneous blockchains, especially in environments with varying trust levels, need to work together to complete a transaction or task. Additionally, scalability remains a challenge, as the computational and storage requirements for managing large-scale value networks can be substantial. As the number of LM agents and transactions grows, ensuring that the underlying blockchain infrastructure can scale to meet demand without compromising performance or security is crucial.
VII Conclusion
In this paper, we have provided an in-depth survey of the state-of-the-art in the architecture, cooperation paradigms, security and privacy, and future trends of LM agent networks. Specifically, we have introduced a novel architecture and its key components, critical characteristics, enabling technologies, and potential applications, toward autonomous, embodied, and connected intelligence of LM agents. Afterward, we have explored the taxonomy of interaction patterns and practical collaboration paradigms among LM agents, including data, computation, and information sharing for collective intelligence. Furthermore, we have identified significant security and privacy threats inherent in the ecosystem of LM agents, discussed the challenges of security/privacy protections in multi-agent environments, and reviewed existing and potential countermeasures. As the field progresses, ongoing research and innovation will be crucial for overcoming existing limitations and harnessing the full potential of LM agents in transforming intelligent systems.